### v1

# Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning

Abstract
Reasoning is a core capability in artificial intelligence systems, for which large language models (LLMs) have recently shown remarkable progress. However, most work focuses exclusively on deductive reasoning, which is problematic since other types of reasoning are also essential in solving real-world problems, and they are less explored. This work focuses on evaluating LLMs’ inductive and abductive reasoning capabilities. We introduce a programmable and synthetic dataset, InAbHyD (pronounced in-a-bid), where each reasoning example consists of an incomplete world model and a set of observations. The task for the intelligent agent is to produce hypotheses to explain observations under the incomplete world model to solve each reasoning example. We propose a new metric to evaluate the quality of hypotheses based on Occam’s Razor. We evaluate and analyze some state-of-the-art LLMs. Our analysis shows that LLMs can perform inductive and abductive reasoning in simple scenarios, but struggle with complex world models and producing high-quality hypotheses, even with popular reasoning-enhancing techniques such as in-context learning and RLVR.  [^1] All data and generation code are freely available at https://github.com/byrantwithyou/inabhyd/.
1 Introduction
Reasoning is a hallmark of human intelligence. Peirce categorizes reasoning into deductive, inductive, and abductive reasoning  (Burch, 2024). Deductive reasoning draws a specific conclusion from a set of premises. For example, given “All humans are moral” and “Socrates is a human”, we can conclude that “Socrates is moral”. Inductive reasoning aims to identify a general principle from multiple individual cases. For example, if we see many cute cats, we may inductively reason that all cats are cute. Abductive reasoning seeks to produce hypotheses to explain observations. For example, if we hear a bird sound outside, we may abductively reason that there is a bird. The goal in both types of reasoning is to find explanations to explain observations (inductive reasoning focuses on inducing a more general rule, while abductive reasoning focuses on finding a ground fact or axiom).
The composition of inductive and abductive reasoning is ubiquitous in real-world scenarios. For example, given (i) Amy has a skin rash and leg pain and has been diagnosed with a new unknown virus. (ii) Sam has a skin rash and leg pain and has been diagnosed with the same new unknown virus. (iii) Tom also has a skin rash and leg pain. We can first induce the rule: The new unknown virus causes skin rashes and leg pain, and then we can abduce that Tom is infected with this new virus.
A fundamental difference between deductive and inductive or abductive reasoning is that deductive reasoning is deterministic. The conclusion in deductive reasoning is deterministically correct if the proof is sound and the premises are correct. However, inductive and abductive reasoning are inherently probabilistic. Technically, there may exist many conclusions, some more probable (high-quality) and some less probable (low-quality). For example, in the above bird sound case, a less probable conclusion is that someone outside is playing the bird sound on a speaker. In this regard, we view inductive and abductive reasoning as the inverse processes of deductive reasoning and study them through a unified lens. In many real-world scenarios, all three reasoning modalities are indispensable. Deductive reasoning is widely used in mathematical problem-solving, inductive reasoning is commonly used in scientific research and discovery, and abductive reasoning is often employed in medical diagnosis. Knowing and applying the principles of parsimony / Occam’s Razor is essential for LLMs to be utilized in areas such as scientific discovery, which require inductive and abductive reasoning, yet this is often overlooked. Physicists, for example, strive for the simplest explanations to account for as many observations as possible. According to our findings, LLMs struggle to generate high-quality hypotheses that match human-level intelligence. To the best of our knowledge, prior studies of inductive and abductive reasoning did not consider parsimony / Occam’s Razor in their evaluation.
The high-level research question we want to answer in this paper is: Do LLMs reason inductively and abductively? If so, how well? There is a significant body of work on evaluating the deductive reasoning capabilities of LLMs  (Saparov and He, 2023; Saparov et al., 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024), as well as techniques to improve deductive reasoning capabilities, such as in-context learning and reinforcement learning with verifiable reward, RLVR  (Lambert et al., 2024; Su et al., 2025). However, there is relatively little exploration into the inductive and abductive reasoning capabilities of large language models (LLMs). This work systematically evaluates whether and to what extent LLMs can reason abductively and inductively. We also assess whether popular reasoning-enhancing techniques, such as in-context learning  (Zhou et al., 2022; Dong et al., 2024) and RLVR  (Lambert et al., 2024; Su et al., 2025), which primarily aim to improve deductive reasoning capabilities, can also benefit inductive and abductive reasoning.
Q: All mammals are warm-blooded. Sam is a mammal. Tom is a mammal. John is a mammal. Each feline is a mammal. All canidae are mammals. All canidae are muscular. Every dog is a canidae. All wolves are canidae. All felines are not slow. Each cat is a feline. All tigers are felines. All cats are cute. Alice is a cat. All dogs are smart. Bob is a dog. All wolves are gregarious. Jessica is a wolf. All tigers are strong. Susan is a tiger. All rodents are gnawing. Each rat is a rodent. Each squirrel is a rodent. All rats are adaptable. Jack is a rat. Noah is a rat. All squirrels are bushy. Oliver is a squirrel. We observe that Sam is hairy. Alice is hairy. Bob is hairy. Fae is strong. Fae is not slow. Fae is warm-blooded. Jack is a mammal. Noah is a mammal. Oliver is a mammal.
Please produce hypotheses to explain all observations.
A: All mammals are hairy. Fae is a tiger. All rodents are mammals.
world model observations hypotheses
Figure 1: An example of a reasoning question in InAbHyD . The upper left side is the world model. Note that the dashed arrow means the subtype relationship All rats are mammals is hidden and needs to be inferred (i.e., a hypothesis to propose). Each node is a concept (e.g., cat), with its properties and members. Each edge denotes a subtype relationship (e.g., All cats are carnivores). On the upper right-hand side are observations in this reasoning example and hypotheses (ground truth answers for this reasoning example). Each hypothesis explains observations of the same color (e.g., the hypothesis All mammals are hairy explains the observations Sam is hairy; Alice is hairy; Bob is hairy). On the lower side is the natural language description of this reasoning example. Note that this example uses a real world model for illustrative purposes. In our dataset, we use fictional world models. A more detailed explanation of this example is given in Appendix A.
We create a programmable and synthetic dataset, InAbHyD (Inductive and Abductive Hypothesis Discovery), to test LLMs’ inductive and abductive reasoning capabilities. Each reasoning example has a fictional incomplete world model in the form of an ontology tree and a set of observations, represented in first-order logic. To answer each reasoning example, LLMs must produce hypotheses to explain observations under the incomplete world model. We generate synthetic examples of fictional world models to mitigate training data contamination. The ontology tree’s height characterizes the difficulty of the world model and the corresponding reasoning example. [^2] See Figure 1 for an example and Appendix A for more explanations for Figure 1 and more examples in Appendix B.
We evaluate our dataset on five widely-used LLMs: Llama3-70B  (Dubey et al., 2024) , Gemma3-27B  (Kamath et al., 2025), DeepSeek-V3  (DeepSeek-AI et al., 2024), GPT-4o  (OpenAI, 2023) , and DeepSeek-R1-Distill-Llama-70B  (DeepSeek-AI et al., 2025). For each LLM, we test whether it can produce valid and high-quality hypotheses. We also test whether popular reasoning-enhancing techniques, such as in-context learning and RLVR, which primarily were meant to improve deductive reasoning, can also benefit inductive and abductive reasoning. Our findings can be summarized as follows:
  1. 1.
    LLMs can perform inductive and abductive reasoning at high accuracy for shallow ontology trees with a single hypothesis.
  2. 2.
    The accuracy drops significantly when the ontology tree structure is complex and there are multiple hypotheses.
  3. 3.
    In some cases, LLMs may produce valid hypotheses (hypotheses which can explain all observations); but fail to produce high-quality hypotheses.
  4. 4.
    In general, we observe that reasoning-enhancing techniques for deductive reasoning only moderately benefits inductive and abductive reasoning.
2 Related Work
Evaluating deductive reasoning capabilities.
There is no consensus on whether LLMs can perform general deductive reasoning or are stochastic parrots (Bender et al., 2021). Researchers have designed different datasets and evaluation methodologies to answer this question  (Saparov et al., 2023; Saparov and He, 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024). LogiQA  (Liu et al., 2020) is a multiple-choice QA dataset to test multiple types of deductive reasoning of language models. However, it only attends to label accuracy (the correctness of the final answer) without considering the correctness of intermediate proof steps. In contrast, PrOntoQA  (Saparov and He, 2023) and PrOntoQA-OOD  (Saparov et al., 2023) evaluate both label accuracy and step-by-step proof correctness. They can synthetically generate proofs with different heights, widths, and deduction rules. Our work is similar to PrOntoQA and ProntoQA-OOD in spirit. We both use a first-order ontology tree as the world model. However, PrOntoQA and PrOntoQA-OOD do not evaluate inductive and abductive reasoning.
Evaluating non-deductive reasoning capabilities.
[…]
Dataset Controls for complexity of world models Natural language description Inductive reasoning Abductive reasoning Evaluation of hypotheses quality Multiple hypotheses
ART (Zhao et al., 2023) non-quantitative
[…]
AbductionRules (Young et al., 2022)
Wino-Why (Zhang et al., 2020)
[…]
UniADILR (Sheng et al., 2025)
Mirage (Li et al., 2025)
Case2Code (Shao et al., 2024)
InAbHyD (ours)
Table 1 summarizes the similarities and differences between our work and others on evaluating non-deductive reasoning capabilities in both the pre- and LLM eras  (Young et al., 2022; Hua et al., 2025; Zhao et al., 2023; Yang et al., 2024; Zhang et al., 2021). Our work differs from others in the following ways. First, we focus on inductive and abductive reasoning and their interactions with deductive reasoning, while other works focus only on a single reasoning modality. Second, our reasoning problem is built upon a fictional world model with natural language descriptions. The fictional world model avoids training data contamination compared with ART  (Zhao et al., 2023), where the problem is directly taken from a web corpus; the natural language descriptions of the reasoning problem render the evaluation more practical compared with symbolic reasoning, such as in InductiveBench  (Hua et al., 2025). Third, our dataset is programmable with highly-controllable difficulty parameters, allowing us to evaluate LLMs’ reasoning capabilities under various scenarios.
Techniques to enhance reasoning.
A large body of work has focused on improving the (deductive) reasoning capabilities of LLMs and related downstream tasks such as mathematics and coding  (Li, 2025; Wei et al., 2022; Yao et al., 2023; Wang et al., 2023; Su et al., 2025; Lambert et al., 2024; Zhou et al., 2022; Dong et al., 2024). The first line of work aims to elicit the emergent capabilities of LLMs by different formats of prompts, such as chain-of-thought (CoT)  (Wei et al., 2022), tree-of-thought (ToT)  (Yao et al., 2023), and self-consistency  (Wang et al., 2023). The second line of work leverages in-context learning by providing few-shot examples to LLMs. The third line of work views LLM token generation as a reinforcement learning problem and utilizes verifiable reward functions (reinforcement learning with verifiable reward, RLVRLambert et al. 2024; Su et al. 2025) to steer LLMs to improved reasoning. The fourth line of work scales LLMs’ test time to increase the accuracy of reasoning problems. It is worth noting that most techniques discussed above focus mainly on deductive reasoning. This work investigates whether the above methods benefit inductive and abductive reasoning.
3 Approach
In this section, we describe how we generate the dataset and evaluate LLMs on this dataset. At a high level, the dataset consists of synthetically-generated reasoning examples in natural language. The question description is a user prompt, and the format specification is a system prompt to LLMs. We parse LLMs’ output to evaluate their inductive and abductive reasoning capabilities on metrics such as accuracy. We acknowledge that the potential space of inductive and abductive reasoning examples is vast, so we restrict ourselves to what is expressible in first-order logic. By composing the inductive and abductive reasoning steps in Table 2, our evaluation framework can readily cover all inductive and abductive reasoning questions in first-order logic.
Each reasoning example contains a fictional world model and a set of observations. We use real person names (such as Amy) and real property names (such as strong), but we do not use real concept names (such as tiger). Instead, we use fictional concept names (such as wumpus). In doing so, LLMs will not be confused by these fictional relations, as we never generate facts that contradict those about real concepts.
LLMs must produce hypotheses to explain the observations. Our hypothesize-explain framework can cover a wide range of inductive/abductive reasoning scenarios. [^3] Traditionally, producing hypotheses to explain observations is considered abductive reasoning. However, we include inductive reasoning in the same framing. For example, given these axioms as a world model: Sally is a cat; Polly is a cat; Amy is a cat and these observations: Sally is cute; Polly is cute; Amy is cute, we can produce a hypothesis: All cats are cute to explain all observations. This is inductive reasoning. To generate a single reasoning example, we go through the following steps:
  1. 1.
    First, we construct a world model as an ontology tree in first-order logic.
  2. 2.
    Second, we traverse the ontology tree and randomly mark some axioms as hidden. Each hidden axiom is the ground truth for hypotheses.
  3. 3.
    Third, for each hidden axiom, we generate observations so that LLMs have enough information to produce ground truth hypotheses.
  4. 4.
    Fourth, we translate the incomplete world model and observations into natural language.
Subtask Reasoning type tested World model Observations Ground truth hypothesis Proof trees for observations
Infer property Inductive o(A)
o(B)
o(C)
p(A)
p(B)
p(C)
[…]
Infer
membership
relation
Abductive $\forall x(o(x)\to p_{1}(x))$
$\forall x(o(x)\to p_{2}(x))$
$\forall x(o(x)\to p_{3}(x))$
p1(A)
p2(A)
p3(A)
[…]
Infer
subtype
relation
Both o1(A)
o1(B)
o1(C)
o2(A)
o2(B)
o2(C)
[…]
Table 2: An overview of various subtasks in InAbHyD and their details in the simplest case (a height-1 ontology tree with a single hypothesis). Here o denotes an ontology concept (e.g, cat) and o(A) denotes the membership (e.g., cat(Sally) means Sally is a cat). p denotes a property (e.g, cute) and p(A) denotes A has property p (e.g., cute(Sally) means Sally is cute). We only show the proof tree for one observation for simplicity; other observations can be derived similarly. We show hypotheses green and axioms in world models purple in the proof tree.
See Table 2 for an overview of the different subtasks and their details in the simplest case (a height-1 ontology tree with a single hypothesis). We evaluate the output of LLMs (i.e., produced hypotheses) on the following metrics:
  1. 1.
    Strong accuracy: Hypotheses are strongly correct if they exactly match ground truth hypotheses.
  2. 2.
    Weak accuracy: Hypotheses are weakly correct if they explain all observations.
  3. 3.
    Hypothesis quality: A quantitative measurement of the quality of the LLMs’ produced hypotheses based on the principles of parsimony.
Next, we describe each step in more detail.
World model generation
In our approach, the world model is an ontology tree in first-order logic. Each node in the ontology tree is a concept (e.g. cat) with properties (e.g. All cats are cute, or equivalently in logical form $\forall x(\texttt{cat}(x)\rightarrow\texttt{cute}(x))$) and members (e.g. Amy is a cat, or equivalently in logical forms, cat(Amy)). Each edge in the ontology tree denotes a subtype relationship (e.g., the ragdoll node being a child of the cat node implies Each ragdoll is a cat, or equivalently in logical form, $\forall x(\texttt{ragdoll}(x)\rightarrow\texttt{cat}(x))$). We first generate the topology of the ontology tree with the height as a controllable parameter. Each non-leaf node has two or three child nodes. The height of the ontology tree is a variable that characterizes the complexity of the world model. We then make two passes of the tree in level order traversal. In the first pass, for each node, we allocate a concept name, three properties, [^4] The properties can also be negated properties, e.g., All cats are not slow. and three members. After the first pass, we have a complete world model with three types of axioms: (1) properties (e.g., All cats are cute), (2) membership relations (e.g., Amy is a cat), (3) subtype relations (e.g., Each ragdoll is a cat). In the second pass, we randomly mark some axioms as hidden. These axioms are the ground truth hypotheses. After the second pass, we have an incomplete world model.
Observation generation
Next, we generate observations so that LLMs have sufficient information to produce the ground truth hypotheses. The observations must be complex enough that the reasoning example is non-trivial, yet they should also ensure the reasoning is feasible.
  • To generate an example where the task is to infer a property (e.g., All cats are cute), we first identify the type in the ontology with this property (e.g., cat), then we sample members of this type and subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where the members have this property (e.g., Amy is cute; Sally is cute; Polly is cute).
  • To generate an example where the task is to infer a membership relation (e.g., Amy is a cat), we first identify the type in the ontology with this member (e.g., cat), then we sample properties of this type and supertypes (e.g., All cats are cute; All felines are not slow; All mammals are hairy), and add observations where this member have these properties (e.g., Amy is cute; Amy is not slow; Amy is hairy).
  • To generate an example where the task is to infer a subtype relation (e.g., All cats are felines), we first identify the subtype in the ontology (e.g., cat), then we sample members of this type and its subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where these members belong to the supertype (e.g., Amy is a feline; Polly is a feline; Sally is a feline).
Translate theories and observations into natural language
We use the same grammar as in PrOntoQA for translating logical forms into natural language. Properties are translated into the form: All cats are cute; subtype relations are randomly translated into one of three forms: Each ragdoll is a cat; Every ragdoll is a cat; All ragdolls are cats; membership relations are translated into the form Amy is a cat.
Hypothesis quality
We define a metric to evaluate the quality of hypotheses based on the principles of parsimony, or Occam’s Razor (i.e., the priciple that, other things being equal, simpler theories are better;  Baker 2022). In this regard, we encourage one hypothesis to explain as many observations as possible and penalize unnecessary hypotheses. Formally, given a world model $\mathcal{T}$, a set of observations $\mathcal{O}$, their associated proof trees $\mathcal{P}$,candidate hypotheses $\mathcal{H}$, and ground truth hypotheses $\mathcal{H}^{*}$, for $h\in\mathcal{H}$, $n(h)$ denotes the number of times $h$ appears in the proof trees of $\mathcal{O}$ (i.e., how many times $h$ is used as a premise of a proof step). We use $|\mathcal{H}|$ to denote the cardinality of $\mathcal{H}$. The quality of $\mathcal{H}$, $q(\mathcal{H})$ is defined by  [^5] We write the equation as the ratio of two averages: the numerator is the average accuracy of the predicted hypothesis (i.e., the average number of times a hypothesis appears in any valid proof), and the denominator is the average number of times a hypothesis appears in the shortest (i.e., most parsimonious) proof.:
[…]
According to the definition, the ground truth hypotheses $\mathcal{H}^{*}$ have a quality score of 1. Consider the example in Figure 1, here $\mathcal{H}^{*}$ is {Fae is a tiger, All mammals are hairy, All rodents are mammals} and $|\mathcal{H}^{*}|=3$. Each $h^{*}\in\mathcal{H}^{*}$ appears three times in the proof trees of $\mathcal{O}$, so $\sum_{h^{*}\in\mathcal{H}^{*}}n(h^{*})=9$. Consider candidate hypotheses $\mathcal{H}$ {Fae is a tiger, All mammals are hairy, All rats are mammals, All squirrels are mammals, All cats are hairy}. $|\mathcal{H}|=5$. Obviously $\mathcal{H}$ can explain $\mathcal{O}$. All rats are mammals appears twice, All squirrels are mammals appears once, and All cats are hairy appear once in the proof trees of $\mathcal{O}$, so $\sum_{h\in\mathcal{H}}n(h)=3+3+2+1+1=10$. In this case, $q(\mathcal{H})=\frac{3\times 10}{5\times 9}=0.67$.
Appendix C contains the complete algorithms for generating reasoning examples and evaluating hypotheses.
4 Results
4.1 Experimental setup
In each experiment configuration, we generate 100 independent and identically distributed reasoning examples. Each reasoning example is pure text, a user prompt with an incomplete world model and observations. We also specify the desired output format in the system prompt. To answer each reasoning example, LLMs must produce hypotheses to explain all observations. We evaluate hypotheses on three metrics: weak accuracy, strong accuracy, and quality. We compute 95% confidence intervals for both accuracy metrics using the Wilson interval (Wilson, 1927). We use the OpenAI API to evaluate GPT-4o and together.ai platform to evaluate Llama3-70B, Gemma3-27B, DeepSeek-V3, and DeepSeek-R1-Distill-Llama-70B.
In our experiments, we vary the following variables: (1) Different types of tasks (i.e., infer property, infer membership relation, or infer subtype relation), (2) Single hypothesis or multiple hypotheses, (3) The height of the ontology tree, (4) Whether the reasoning example contains in-context demonstrations and whether they are in-distribution or out-of-distribution.
4.2 Zero-shot results
To begin evaluating the inductive and abductive reasoning abilities of LLMs, we first investigate the simplest setting with zero-shot prompting.
4.2.1 Single hypothesis
We first investigate results for reasoning examples with one hypothesis for three tasks. This helps us to isolate different types of tasks and understand how LLMs perform in each case. The results are in Figure 2. In general, all models show high accuracy (both weak and strong) greater than 80% for all three tasks under a height-1 ontology tree. When the height of the ontology tree starts to increase, we also observe a general decreasing trend in all three metrics, which is not surprising. One interesting observation is that LLMs perform better in infer membership relation tasks than other tasks, especially for Gemma3-27B. For example, even under a height-4 ontology tree, Gemma3-27B can achieve near 50% weak accuracy, while the number drops to around 10% for infer subtype relation tasks. This may be because the weak correctness for infer membership relation tasks only requires searching for the concept that has some property, but do not need to understand the structure of the ontology tree.666e.g., given Amy is cute, LLMs only need to search for a concept which has the property cute to infer the membership relation of Amy. However, this is not the case for infer subtype relation and infer properties tasks. We observe a similar trend for infer subtype relation and infer properties tasks, as we can view the concept’s parent as a special property of that concept.
1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00QualityTask: infer membership relationTask: infer propertyTask: infer subtype relationGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
Figure 2: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with single hypothesis.
4.2.2 Multiple hypotheses
Next, we turn to multi-hypothesis scenarios because they are closer to real-life situations, and we are interested in understanding how LLMs perform in this setting. Each reasoning example can contain multiple hypotheses. To check the difficulty of each reasoning example, we sample the ground truth hypotheses such that the number of hypotheses grows linearly w.r.t. the height of the ontology tree. [^7] Each ground truth hypothesis has $\sim 3$ observations. So the number of observations also grows linearly w.r.t. the height of the ontology tree. Invariably, the number of axioms in the world model grows exponentially w.r.t. the height of the ontology tree. See Table 3 for statistics under multi-hypothesis scenarios and the top row of Figure 3 for results.
The height of the ontology tree # of world model axioms # of observations # of ground truth hypotheses 1 9.0 10.0 3.0 2 14.0 11.8 3.5 3 25.5 15.2 4.6 4 46.8 20.0 6.6
Table 3: Basic statistics of multi-hypothesis examples in InAbHyD . We list the average number of world model axioms, observations, and ground truth hypotheses w.r.t. the height of the ontology tree in the same set of reasoning examples we used in evaluation. Note that the number of the ground truth hypotheses and the number of observations grow linearly with the height of the ontology tree. In contrast, the number of world model axioms grows exponentially.
In general, the performance in the multi-hypothesis setting differed significantly from that in the single-hypothesis setting. For example, the accuracy of all models drops significantly when the height of the ontology tree is increased from 1 to 2, with accuracy going from above 80% to below 50%, except for GPT-4o. This is surprising because the average number of ground truth hypotheses only increases from $3.0$ to $3.5$. We attribute this to the coupling effect of multiple hypotheses to produce and the topology of the ontology tree. Reasoning examples under the ontology tree with height 3 or 4 achieve lower accuracies and quality scores for all models. We also observe a notable gap between weak accuracy and quality. For example, under the height-4 ontology tree, all models can achieve at least 20% weak accuracy, but struggle with extremely low quality and strong accuracy scores. We manually inspect some incorrect hypotheses and summarize the most common types of errors in Section 4.5.
4.3 Does in-context learning help?
[…]
Figure 3: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-short (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom). We list the absolute results of the 8-shot in-distribution and out-of-distribution in Appendix D.
LLMs have shown extraordinary generalization capabilities with few-shot examples coupled with CoT prompts (Brown et al., 2020). We are also interested in whether in-context learning can help with inductive and abductive reasoning. We add eight demonstrations to each reasoning example, with the demonstrations’ ground truth hypotheses and CoTs. The CoT consists of each observation’s step-by-step proof. We have two settings for in-context learning: (1) In-distribution demonstrations, where the demonstration has an ontology tree of the same height as the test question. (2) Out-of-distribution demonstrations, where each demonstration has a single hypothesis with a height-1 ontology tree. In both settings, we ensure demonstrations cover all types of tasks.
The in-distribution demonstrations are slightly more helpful than out-of-distribution demonstrations. This result is consistent with the hypothesis that in-context learning mimics Bayesian inference (Xie et al., 2022). We also observe that in-context learning with in-distribution demonstrations improves strong accuracy and quality in ontology trees with heights 3 and 4. We conjecture that LLMs can learn how to navigate the ontology hierarchy and produce high-quality hypotheses through in-context demonstrations. In contrast, we do not observe a significant performance improvement from out-of-distribution demonstrations.
4.4 Does RLVR help?
1234Height0.00.20.40.60.81.0Weak Accuracy1234Height0.00.20.40.60.81.0Strong Accuracy1234Height0.00.20.40.60.81.0QualityLlama3-70BDeepSeek-R1-Distill-Llama-70B
Figure 4: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree comparison results for Llama3-70B and DeepSeek-R1-Distill-Llama-70B with multiple hypotheses.
RLVR has demonstrated significant improvements in other reasoning benchmarks, and so we test whether this improvement generalizes to abductive and inductive reasoning. Specifically, we compare Llama3-70B and DeepSeek-R1-Distill-Llama-70B. In general, we observe an average performance improvement between 10% and 20%. To analyze why DeepSeek-R1-Distill-Llama-70B performs better than Llama3-70B, we also manually inspect the reasoning output of DeepSeek-R1-Distill-Llama-70B (i.e., tokens that are inside <think> and </think>). After producing hypotheses, DeepSeek-R1-Distill-Llama-70B will usually attempt to verify those hypotheses. This verification process is essentially a deductive reasoning process, and can catch some reasoning errors or fix low-quality hypotheses. A complete example is in Appendix E.
4.5 Analysis of common errors
To further understand how and why LLMs fail, we manually inspect mistaken and low-quality hypotheses. We summarize some of the most common types of errors here for further insights. We provide example responses in each category in Appendix E. (1) Wrong ontology direction.LLMs sometimes produce hypotheses in the wrong direction. For example, the ground truth hypothesis is All cats are mammals, while LLMs will produce the hypothesis All mammals are cats. (2) Ignore the ontology and produce unnecessary hypotheses.LLMs sometimes can produce unnecessary hypotheses because they ignore the ontology. For example, the ground truth hypothesis is All cats are cute, while LLMs produce two hypotheses: All cats are cute; All ragdolls are cute. (3) Fall back to trivial hypotheses.Technically, the observation is also a valid hypothesis, but it is considered low-quality because it can only explain itself. LLMs sometimes reuse observations as hypotheses. (4) Hallucinated entities.LLMs sometimes produce hypotheses with hallucinated entities (including concepts, properties, members). Usually, this happens when the entity appears in the in-context demonstrations, but LLMs use them in the test question. (5) Confuse members with concepts.We use fictional names for concepts and real-person names for members. However, LLMs sometimes do confuse these two. For example, the ground truth hypothesis is Amy is a wumpus, while LLMs produce Wumpus is Amy.
5 Conclusions and future work
In this study, we introduced a programmable and synthetic dataset, InAbHyD . Our dataset serves as an out-of-distribution evaluation of the inductive and abductive reasoning abilities of LLMs. We do not claim that success in our benchmark automatically implies success in real-world tasks, but rather, our benchmark serves as a checkpoint for developing models that are capable of robust real-world abductive and inductive reasoning.
We systematically studied LLMs’ inductive and abductive reasoning capabilities. Our results suggest that current LLMs can’t generate high-quality hypotheses that match human-level intelligence. For example, suppose we use LLMs as a scientific discovery assistant. In that case, they may generate superficial hypotheses that can explain observations but fail to uncover more general hypotheses that reveal hidden structures or connections in science. Our results can serve to better inform practitioners of LLMs of their limitations, especially in applications that rely heavily on inductive and abductive reasoning. The dataset itself opens new avenues for future work, such as training or fine-tuning models to improve their inductive and abductive reasoning capabilities.
One limitation of this work is that first-order logic (FOL) limits the expressiveness of the world model and the reasoning example. We leave it as a future work to use higher-order logic (HOL) to generate more complex world models and reasoning examples, which would require translating from HOL into natural language and vice versa. The other crucial future direction is improving the quality of LLMs’ hypotheses. For that, LLMs may benefit from RLVR  (Lambert et al., 2024) that explicitly conveys principles of parsimony in inductive and abductive reasoning with an appropriate reward. Another interesting future direction will be to determine whether large language models fine-tuned on this synthetic dataset (using supervised fine-tuning or reinforcement learning from verifiable rewards) could improve their performance on real-world tasks involving inductive and abductive reasoning. If this is the case, this synthetic dataset is more valuable because we can essentially generate an infinite number of synthetic data points, without worrying about the bottleneck of sampling data from the real world.
Appendix A Detailed explanations of Figure 1
Here we list proof trees for each observation in Figure 1. Hypotheses are in green and world model axioms are in purple.
Fae is strong
[…]
Fae is not slow
$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\neg\texttt{slow}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\neg\texttt{slow}(x))$     $\texttt{tiger}(\texttt{Fae})$                             $\neg\texttt{slow}(\texttt{Fae})$
Fae is warm-blooded
$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\texttt{mammal}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\texttt{mammal}(x))$     $\forall x(\texttt{mammal}(x)\rightarrow\texttt{warm-blooded}(x))$                              $\forall x(\texttt{tiger}(x)\to\texttt{warm-blooded}(x))$     tiger(Fae)                                                     warm-blooded(Fae)
Sam is hairy
$\forall x(\texttt{mammal}(x)\to\texttt{hairy}(x))$     mammal(Sam)           hairy(Sam)
Alice is hairy
$\forall x(\texttt{cat}(x)\to\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\to\texttt{mammal}(x))$           $\forall x(\texttt{cat}(x)\to\texttt{mammal}(x))$     cat(Alice)                           mammal(Alice)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Alice)
Bob is hairy
$\forall x(\texttt{dog}(x)\to\texttt{canidae}(x))$     $\forall x(\texttt{canidae}\to\texttt{mammal}(x))$            $\forall x(\texttt{dog}(x)\to\texttt{mammal}(x))$     dog(Bob)                             mammal(Bob)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Bob)
Jack is a mammal
rat(Jack)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Jack)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Jack)
Noah is a mammal
rat(Noah)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Noah)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Noah)
Oliver is a mammal
squirrel(Oliver)     $\forall x(\texttt{squirrel}(x)\to\texttt{rodent}(x))$             rodent(Oliver)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                                  mammal(Oliver)
Appendix B More examples
Single hypothesis; task: infer property; height-1 ontology tree.
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
Hypotheses: Dalpists are rainy.
Single hypothesis; task: infer property; height-2 ontology tree.
World Model: Rompuses are dalpists. Sarpers are dalpists. Each gergit is a dalpist. Pamela is a sarper. Jerry is a dalpist. Amy is a rompus.
Observations: Amy is rainy. Pamela is rainy. Jerry is rainy.
Hypotheses: Dalpists are rainy.
Single hypothesis; task: infer property; height-3 ontology tree.
World Model: Pamela is a porpor. Jerry is a dalpist. Twimpees are sarpers. Each gergit is a dalpist. Each storpist is a gergit. Amy is a storpist. Porpors are rompuses. Sarpers are dalpists. Rompuses are dalpists.
Observations: Jerry is not muffled. Amy is not muffled. Pamela is not muffled.
Hypotheses: Each dalpist is not muffled.
Single hypothesis; task: infer property; height-4 ontology tree.
World Model: Amy is a felper. Jerry is a dalpist. Pamela is a felper. Each rompus is a dalpist. Each porpor is a rompus. Sarpers are dalpist s. Shumples are twimpees. Gergits are dalpists. Each frompor is a porpor. Storpists are gergits. Twimpees are sarpers. Felpers are storpists.
Observations: Amy is translucent. Pamela is translucent. Jerry is translucent.
Hypotheses: Each dalpist is translucent.
Single hypothesis; task: infer membership relation; height-1 ontology tree.
World Model: Dalpists are liquid. Dalpists are rainy. Each dalpist is brown.
Observations: Jerry is brown. Jerry is rainy. Jerry is liquid.
Hypotheses: Jerry is a dalpist.
Single hypothesis; task: infer membership relation; height-2 ontology tree.
World Model: Rompuses are dalpists. Jerry is a dalpist. Each sarper is a dalpist. Each dalpist is brown. Edward is a sarper. Gergits are dalpists. Gergits are rainy. Amy is a rompus. Dalpists are liquid.
Observations: Pamela is rainy. Pamela is liquid. Pamela is brown.
Hypotheses: Pamela is a gergit.
Single hypothesis; task: infer membership relation; height-3 ontology tree.
World Model: Each sarper is a dalpist. Jerry is a dalpist. Barbara is a porpor. Each dalpist is moderate. Rompuses are dalpists. Twimpees are sarpers. Storpists are gergits. Each dalpist is large. Edward is a sarper. Pamela is a gergit. Gergits are dalpists. Debra is a storpist. Porpors are rompuses. Amy is a rompus. Twimpees are not muffled.
Observations: Nicole is moderate. Nicole is large. Nicole is not muffled.
Hypotheses: Nicole is a twimpee.
Single hypothesis; task: infer membership relation; height-4 ontology tree.
World Model: Amy is a rompus. Each porpor is a rompus. Each sarper is a dalpist. Sharon is a felper. Rompuses are dalpists. Each shumple is a twimpee. Each dalpist is fruity. Barbara is a porpor. Every frompor is a porpor. Every gergit is a dalpist. Edward is a sarper. Raymond is a frompor. Each storpist is a gergit. Jerry is a dalpist. Each dalpist is moderate. Shumples are translucent. Felpers are storpists. Nicole is a twimpee. Pamela is a gergit. Twimpees are sarpers. Debra is a storpist.
Observations: Michael is fruity. Michael is translucent. Michael is moderate.
Hypotheses: Michael is a shumple.
Single hypothesis; task: infer subtype relation; height-1 ontology tree.
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is a rompus. Jerry is a rompus. Pamela is a rompus.
Hypotheses: Dalpists are rompuses.
Single hypothesis; task: infer subtype relation; height-2 ontology tree.
World Model: Gergits are dalpists. Porpors are dalpists. Each sarper is a dalpist. Pamela is a porpor. Jerry is a dalpist. Amy is a gergit.
Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
Hypotheses: Dalpists are rompuses.
Single hypothesis; task: infer subtype relation; height-3 ontology tree.
World Model: Amy is a storpist. Jerry is a dalpist. Frompors are porpors. Each sarper is a dalpist. Each twimpee is a sarper. Pamela is a twimpee. Storpists are gergits. Porpors are dalpists. Gergits are dalpists.
Observations: Jerry is a rompus. Amy is a rompus. Pamela is a rompus.
Hypotheses: Each dalpist is a rompus.
Single hypothesis; task: infer subtype relation; height-4 ontology tree.
World Model: Each sarper is a dalpist. Jerry is a dalpist. Amy is a shumple. Pamela is a shumple. Each storpist is a gergit. Dumpuses are frompors. Gergits are dalpists. Felpers are storpists. Each twimpee is a sarper. Frompors are porpors. Shumples are twimpees. Porpors are dalpists.
Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
Hypotheses: Each dalpist is a rompus.
Multiple hypothesis; height-1 ontology tree.
World Model: Debra is a dalpist. Amy is a dalpist. Pamela is a dalpist. Edward is a dalpist. Dalpists are liquid. Each dalpist is moderate. Each dalpist is brown. Barbara is a dalpist.
Observations: Barbara is rainy. Jerry is a rompus. Jerry is liquid. Jerry is brown. Pamela is a rompus. Edward is rainy. Debra is rainy. Jerry is moderate. Amy is a rompus.
Hypotheses: Dalpists are rompuses. Dalpists are rainy. Jerry is a dalpist.
Multiple hypothesis; height-2 ontology tree.
World Model: Nicole is a gergit. Porpors are dalpists. Pamela is a sarper. Sharon is a gergit. Porpors are not cold. Dalpists are rainy. Raymond is a gergit. Each porpor is translucent. Each sarper is a dalpist. Jerry is a dalpist. Edward is a dalpist. Every gergit is a dalpist. Debra is a dalpist. Barbara is a sarper.
Observations: Barbara is liquid. Amy is translucent. Sharon is moderate. Debra is liquid. Amy is rainy. Raymond is moderate. Jerry is a rompus. Amy is not cold. Pamela is a rompus. Nicole is moderate. Edward is liquid. Amy is a rompus.
Hypotheses: Dalpists are rompuses. Dalpists are liquid. Every gergit is moderate. Amy is a porpor.
Multiple hypothesis; height-3 ontology tree.
World Model: Barbara is a sarper. Sarpers are dalpists. Debra is a shumple. Every felper is a sarper. Raymond is a porpor. Sharon is a storpist. Each storpist is a gergit. Helen is a frompor. Porpors are dalpists. Each dalpist is liquid. Amy is a sarper. Each dalpist is salty. Felpers are rainy. Every shumple is a porpor. Each twimpee is a sarper. Nicole is a gergit. Each frompor is a sarper. Jerry is a dalpist. Dalpists are not muffled. Each gergit is a dalpist. Michael is a twimpee. Sarpers are amenable.
Observations: Debra is large. Amy is a rompus. Pamela is a rompus. Edward is liquid. Pamela is rainy. Edward is not muffled. Barbara is large. Pamela is not muffled. Edward is large. Edward is salty. Jerry is a rompus. Pamela is amenable.
Hypotheses: Each dalpist is a rompus. Dalpists are large. Edward is a dalpist. Pamela is a felper.
Multiple hypothesis; height-4 ontology tree.
World Model: Every kurpor is a twimpee. Each gergit is pale. Patricia is a dumpus. Timples are frompors. Sharon is a folpee. George is an orgit. Rimpees are felpers. Stephen is a shumple. Mark is an orgit. Joseph is a porpor. Andrew is an orgit. Each tumpus is a felper. Gergits are dalpists. Felpers are not dark. Helen is a timple. Jason is a tumpus. Amy is a pergit. Rimpees are angry. Porpors are dalpists. Raymond is a twimpee. Each rimpee is moderate. Linda is a tumpus. Every shumple is a porpor. Each frompor is a sarper. Every dumpus is a storpist. Gergits are amenable. Barbara is an orgit. Rachel is a frompor. Felpers are sarpers. Jacob is a felper. Janet is a tumpus. Michael is a timple. Steven is a boompist. Edward is a dalpist. Storpists are gergits. Each boompist is a frompor. Debra is a dalpist. Nicole is a sarper. Pergits are felpers. Shirley is a timple. Jerry is a dalpist. Pame la is a kurpor. Every folpee is a frompor. Each sarper is a dalpist. Eric is a storpist. Orgits are shumples. Gergits are not o paque. Twimpees are sarpers.
Observations: Linda is brown. Helen is salty. George is translucent. Jason is brown. Raymond is large. Sharon is large. Jerry is a rompus. William is angry. Barbara is liquid. Edward is liquid. Emily is amenable. Janet is brown. Emily is not opaque. Shirley is salty. William is not dark. Pamela is a rompus. Nicole is large. William is moderate. Andrew is translucent. Michael is salty. Amy is a rompus. Debra is liquid. Emily is pale. Mark is translucent.
Hypotheses: Dalpists are rompuses. Every dalpist is liquid. Emily is a gergit. Sarpers are large. Timples are salty. William is a rimpee. Tumpuses are brown. Orgits are translucent
Appendix C Algorithms
Refer to Algorithm 1 and 2.
if task is infer property then
 select a member of the root node
 randomly select a member of the leaf nodes
 randomly select a member of nonleaf nodes
 generates the observations where selected members have the property
end if
if task is infer membership then
 select a random leaf node
 randomly select a property of the leaf node
 randomly select a property of the root node
 randomly select a property of the selected node’s non-root ancestors.
 generates the observations where this entity have all properties
end if
if task is infer subtype relation then
 select a member of the root node
 randomly select a member of the leaf nodes
 randomly select a member of nonleaf nodes
 generates the observations where selected members also belong to a supertype
end if
return all world model axioms, observations, and the ground truth hypothesis
Algorithm 1 Generate a single hypothesis
// First, build the ontology tree
// $Node()$ will initialize a node object
// Add a root node
$tree\_nodes\leftarrow[[Node()]]$
for $i$ from 1 to h1h-1 do
$current\_nodes\leftarrow[]$
for $parrent\_node\in tree\_nodes[-1]$ do
  $u\leftarrow\texttt{random}(0,1)$
  if $u>0.5$ then
   for $j$ from 1 to 2 do
    current_nodes.append(Node())current\_nodes.append(Node())
   end for
  else
   for $j$ from 1 to 3 do
    current_nodes.append(Node())current\_nodes.append(Node())
   end for
  end if
end for
tree_nodes.append(current_nodes)tree\_nodes.append(current\_nodes)
end for
for $current\_nodes\in tree\_nodes$ do
for $node\in current\_nodes$ do
  // Allocate three members and properties
  for $j$ from 1 to 3 do
   $node.members.append(RandomMember())$
   $node.properties.append(RandomProperty())$
  end for
end for
end for
if single hypothesis then
 Randomly select a task type
 generates a single hypothesis
else
 generate 1 to 3 hypothesis per tree layer
1end if
Algorithm 2 Generate a reason example
Appendix D More results
Here we list the absolute results of Figure 3 in Figure 5.
1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality8-shot, out-of-distribution demonstration8-shot, in-distribution demonstrationzero-shotGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
Figure 5: Absolute results of weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-shot (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom).
Appendix E LLM example responses
Self-verification
World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
Hypotheses: Dalpists are rainy.
LLM Output: <think> … Suppose Dalpists are rainy. Since Amy and Jerry and Pamela are dalpists, they are all rainy. This hypothesis can explain all observations. </think>
Wrong ontology direction
World Model: Porpors are dalpists. Nicole is a gergit. Sarpers are dalpists. Every gergit is a dalpist. Debra is a dalpist. Each dalpist is rainy. Every porpor is not cold. Raymond is a gergit. Each porpor is translucent. Amy is a porpor. Jerry is a dalpist. Pamela is a dalpist. Sharon is a gergit. Edward is a dalpist. Michael is a sarper.
Observations: Sharon is moderate. Barbara is not cold. Debra is liquid. Barbara is rainy. Barbara is liquid. Barbara is translucent. Amy is a rompus. Raymond is moderate. Jerry is a rompus. Edward is liquid. Pamela is a rompus. Nicole is moderate.
Ground truth hypotheses: Dalpists are rompuses. Every dalpist is liquid. Each gergit is moderate. Barbara is a porpor.
LLM Output: … Rompuses are dalpists. …
Ignore the ontology and produce unnecessary hypotheses
World Model: David is a sorple. Each sorple is a dropant. Each hilper is a thorpin. Rimpees are scrompists. Brian is a sorple. Dropants are gompers. Angela is a gomper. Each thorpin is a gomper. Every scrompist is a gomper.
Observations: David is rainy. Brian is rainy. Angela is rainy.
Ground truth hypotheses: Every gomper is rainy.
LLM Output: … Every gomper is rainy. Every dropant is rainy. …
Fall back to rival hypothesee
World Model: Rorpants are dolpers. Every worple is a wumpus. Each drompee is moderate. Ashley is a rorpant. Fimples are dolpers. Ronald is a wumpus. Wumpuses are gwompants. Yompins are dolpers. Brian is a dolper. Michelle is a worple. Each dolper is a gwompant. Shirley is a rorpant. Sandra is a bempin. Linda is a fimple. Drompees are pale. Jack is a yompin. Each drompee is a bempin. Bempins are cold. Bempins are gwompants. Emma is a gwompant. Michael is a gwompant.
Observations: Shirley is a borpin. Christopher is pale. Emma is a borpin. Brian is salty. Michael is salty. Christopher is a borpin. Christopher is cold. Ashley is salty. Christopher is moderate.
Ground truth hypotheses: Each gwompant is a borpin. Each gwompant is salty. Christopher is a drompee.
LLM Output: … Shirley is a borpin. …
Hallucinated entities
World Model: Jonathan is a rifpist. Yimples are dull. Samantha is a rifpist. Karen is a gwompant. Scott is a jelgit. Olivia is a frompor. Maria is a lirpin. Twimpees are lirpins. Each rifpist is aggressive. Frompors are jelgits. Jessica is a jelgit. Amanda is a twimpee. Rifpists are jelgits. Each lirpin is a jelgit. Yimples are small. Each gwompant is a frompor. Each yimple is a rifpist.
Observations: Mary is not overcast. Scott is not overcast. Jonathan is not overcast. Samantha is a kurpor. Mary is aggressive. Jessica is a kurpor. Olivia is a kurpor. Mary is small. Mary is dull.
Ground truth hypotheses: Every jelgit is a kurpor. Jelgits are not overcast. Mary is a yimple.
LLM Output: … Mary is a dolphant. …
Confuse members with concepts
World Model: Delpees are impuses. Yumpuses are delpees. Rebecca is an impus. Kathleen is a rifpist. Each impus is rainy. Thomas is a delpee. Lompees are arpers. Each bempin is an impus. Stephen is a bempin. Every impus is shy. Laura is a yumpus. Every arper is an impus. Stephanie is an arper. Lompees are moderate. Rifpists are bempins.
Observations: Frank is moderate. Frank is rainy. Frank is shy.
Ground truth hypotheses: Frank is a lompee.
LLM Output: … Lompee is Frank. …

### v2

@@
-Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning
+Do Language Models Follow Occam’s Razor? An Evaluation of Parsimony in Inductive and Abductive Reasoning
 Abstract
-Reasoning is a core capability in artificial intelligence systems, for which large language models (LLMs) have recently shown remarkable progress.
-However, most work focuses exclusively on deductive reasoning, which is problematic since other types of reasoning are also essential in solving real-world problems, and they are less explored.
-This work focuses on evaluating LLMs’ inductive and abductive reasoning capabilities.
-We introduce a programmable and synthetic dataset, InAbHyD (pronounced in-a-bid), where each reasoning example consists of an incomplete world model and a set of observations.
-The task for the intelligent agent is to produce hypotheses to explain observations under the incomplete world model to solve each reasoning example.
-We propose a new metric to evaluate the quality of hypotheses based on Occam’s Razor.
-We evaluate and analyze some state-of-the-art LLMs.
-Our analysis shows that LLMs can perform inductive and abductive reasoning in simple scenarios, but struggle with complex world models and producing high-quality hypotheses, even with popular reasoning-enhancing techniques such as in-context learning and RLVR.
-[^1] All data and generation code are freely available at https://github.com/byrantwithyou/inabhyd/.
+Non-deductive reasoning, encompassing inductive and abductive reasoning, is essential in addressing complex real-world questions.
+One key feature of inductive and abductive reasoning is that there are many valid hypotheses; the simplest ones (those that adhere to Occam’s Razor) are often most useful.
+However, this aspect is ignored in recent work that evaluates the non-deductive reasoning capabilities of large language models (LLMs).
+This work fills this gap, focusing on understanding whether the inductive and abductive reasoning capabilities of LLMs adhere to Occam’s Razor, while also examining the correctness of their reasoning.
+To accomplish this goal, we introduce a framework to synthetically generate reasoning questions that (a) require inductive reasoning and abductive reasoning simultaneously; (b) is readily extended to produce any abductive/inductive reasoning question expressible in first-order logic.
+The task for the intelligent agent is to produce hypotheses to explain observations under a given world model.
+We also propose a new automated metric to assess whether hypotheses quantitatively adhere to Occam’s Razor; those hypotheses that are correct and simplest are considered high-quality.
+Our findings on state-of-the-art LLMs suggest that LLMs can perform inductive and abductive reasoning in simple scenarios, but struggle with complex world models and with producing high-quality hypotheses, even with popular reasoning-enhancing techniques such as in-context learning and RLVR.
+[…]
+[…]
+Do Language Models Follow Occam’s Razor?
+An Evaluation of Parsimony in Inductive and Abductive Reasoning
+Yunxin Sun  and Abulhair Saparov {sun1114, asaparov}@purdue.edu Purdue University
 1 Introduction
-Reasoning is a hallmark of human intelligence.
-Peirce categorizes reasoning into deductive, inductive, and abductive reasoning  (Burch, 2024).
-Deductive reasoning draws a specific conclusion from a set of premises.
-For example, given “All humans are moral” and “Socrates is a human”, we can conclude that “Socrates is moral”.
+Reasoning is a hallmark of intelligence.
+Peirce categorizes reasoning into deductive, inductive, and abductive reasoning  (peirce1878deduction; cat), all of which are indispensable in real-world scenarios.
+Deductive reasoning draws a conclusion from a set of premises.
 Inductive reasoning aims to identify a general principle from multiple individual cases.
-For example, if we see many cute cats, we may inductively reason that all cats are cute.
 Abductive reasoning seeks to produce hypotheses to explain observations.
-For example, if we hear a bird sound outside, we may abductively reason that there is a bird.
-The goal in both types of reasoning is to find explanations to explain observations (inductive reasoning focuses on inducing a more general rule, while abductive reasoning focuses on finding a ground fact or axiom).
-The composition of inductive and abductive reasoning is ubiquitous in real-world scenarios.
-For example, given (i) Amy has a skin rash and leg pain and has been diagnosed with a new unknown virus.
-(ii) Sam has a skin rash and leg pain and has been diagnosed with the same new unknown virus.
-(iii) Tom also has a skin rash and leg pain.
-We can first induce the rule: The new unknown virus causes skin rashes and leg pain, and then we can abduce that Tom is infected with this new virus.
-A fundamental difference between deductive and inductive or abductive reasoning is that deductive reasoning is deterministic.
-The conclusion in deductive reasoning is deterministically correct if the proof is sound and the premises are correct.
-However, inductive and abductive reasoning are inherently probabilistic.
-Technically, there may exist many conclusions, some more probable (high-quality) and some less probable (low-quality).
-For example, in the above bird sound case, a less probable conclusion is that someone outside is playing the bird sound on a speaker.
-In this regard, we view inductive and abductive reasoning as the inverse processes of deductive reasoning and study them through a unified lens.
-In many real-world scenarios, all three reasoning modalities are indispensable.
-Deductive reasoning is widely used in mathematical problem-solving, inductive reasoning is commonly used in scientific research and discovery, and abductive reasoning is often employed in medical diagnosis.
-Knowing and applying the principles of parsimony / Occam’s Razor is essential for LLMs to be utilized in areas such as scientific discovery, which require inductive and abductive reasoning, yet this is often overlooked.
-Physicists, for example, strive for the simplest explanations to account for as many observations as possible.
-According to our findings, LLMs struggle to generate high-quality hypotheses that match human-level intelligence.
-To the best of our knowledge, prior studies of inductive and abductive reasoning did not consider parsimony / Occam’s Razor in their evaluation.
-The high-level research question we want to answer in this paper is: Do LLMs reason inductively and abductively?
-If so, how well?
-There is a significant body of work on evaluating the deductive reasoning capabilities of LLMs  (Saparov and He, 2023; Saparov et al., 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024), as well as techniques to improve deductive reasoning capabilities, such as in-context learning and reinforcement learning with verifiable reward, RLVR  (Lambert et al., 2024; Su et al., 2025).
-However, there is relatively little exploration into the inductive and abductive reasoning capabilities of large language models (LLMs).
-This work systematically evaluates whether and to what extent LLMs can reason abductively and inductively.
-We also assess whether popular reasoning-enhancing techniques, such as in-context learning  (Zhou et al., 2022; Dong et al., 2024) and RLVR  (Lambert et al., 2024; Su et al., 2025), which primarily aim to improve deductive reasoning capabilities, can also benefit inductive and abductive reasoning.
+See Appendix A for examples of different types of reasoning and for further background.
 Q: All mammals are warm-blooded.
@@
 world model observations hypotheses
-Figure 1: An example of a reasoning question in InAbHyD .
-The upper left side is the world model.
+Figure 1: An example of a question requiring inductive and abductive reasoning.
 Note that the dashed arrow means the subtype relationship All rats are mammals is hidden and needs to be inferred (i.e., a hypothesis to propose).
-Each node is a concept (e.g., cat), with its properties and members.
-Each edge denotes a subtype relationship (e.g., All cats are carnivores).
-On the upper right-hand side are observations in this reasoning example and hypotheses (ground truth answers for this reasoning example).
-Each hypothesis explains observations of the same color (e.g., the hypothesis All mammals are hairy explains the observations Sam is hairy; Alice is hairy; Bob is hairy).
-On the lower side is the natural language description of this reasoning example.
-Note that this example uses a real world model for illustrative purposes.
+This example uses a real-world model for illustrative purposes.
 In our dataset, we use fictional world models.
-A more detailed explanation of this example is given in Appendix A.
-We create a programmable and synthetic dataset, InAbHyD (Inductive and Abductive Hypothesis Discovery), to test LLMs’ inductive and abductive reasoning capabilities.
+There is a significant body of work on evaluating the deductive reasoning capabilities of LLMs  (prontoqa; prontoood; logiqa; proofwriter; follo), as well as techniques to improve deductive reasoning capabilities, such as in-context learning and reinforcement learning with verifiable rewards, RLVR  (tulu; RLVRmultiple).
+However, there is relatively little exploration into the inductive and abductive reasoning capabilities of large language models (LLMs).
+A fundamental difference between deductive and inductive or abductive reasoning is that deductive reasoning is deterministic.
+However, inductive and abductive reasoning are inherently probabilistic.
+There may exist many hypotheses, some more probable and some less probable.
+For example, if we hear a bird sound, a more probable hypothesis is that there is a bird outside; while a less probable hypothesis is that someone outside is playing the bird sound on a speaker.
+Among these possible hypotheses, the simplest ones are often more useful, and we view them as more probable.
+This principle is known as Occam’s Razor.
+We call hypotheses that adhere to Occam’s Razor high-quality hypotheses.
+Knowing and applying Occam’s Razor is essential in areas such as scientific discovery, which require inductive and abductive reasoning, yet this is often overlooked.
+Physicists, for example, strive for the simplest explanations to account for as many observations as possible.
+The high-level research question we aim to answer is: Do LLMs reason inductively and abductively?
+If so, how well do they adhere to Occam’s Razor?
+This work systematically evaluates whether and to what extent LLMs can reason inductively and abductively and adhere to Occam’s Razor.
+We also assess whether popular reasoning-enhancing techniques, such as in-context learning  (ICL1; ICL2) and RLVR  (tulu; RLVRmultiple), which primarily aim to improve deductive reasoning capabilities, can also improve inductive and abductive reasoning.
+To answer this research question, we introduce a framework to synthetically generate reasoning questions that (a) contain inductive reasoning and abductive reasoning simultaneously; (b) is readily extended to produce any inductive/abductive reasoning question expressible in first-order logic.
 Each reasoning example has a fictional incomplete world model in the form of an ontology tree and a set of observations, represented in first-order logic.
 To answer each reasoning example, LLMs must produce hypotheses to explain observations under the incomplete world model.
-We generate synthetic examples of fictional world models to mitigate training data contamination.
+We generate synthetic examples of fictional world models to mitigate the contamination of training data.
 The ontology tree’s height characterizes the difficulty of the world model and the corresponding reasoning example.
-[^2] See Figure 1 for an example and Appendix A for more explanations for Figure 1 and more examples in Appendix B.
-We evaluate our dataset on five widely-used LLMs: Llama3-70B  (Dubey et al., 2024) , Gemma3-27B  (Kamath et al., 2025), DeepSeek-V3  (DeepSeek-AI et al., 2024), GPT-4o  (OpenAI, 2023) , and DeepSeek-R1-Distill-Llama-70B  (DeepSeek-AI et al., 2025).
+[^1] Refer to Figure 1 and Appendix C for examples.
+To assess whether LLMs adhere to Occam’s Razor quantitatively, we also introduce an automated metric to quantify the quality of each hypothesis (i.e., how well each hypothesis adheres to Occam’s Razor).
+We also introduce a new benchmark dataset, InAbHyD(Inductive and Abductive Hypothesis Discovery), generated under this framework.
+This dataset contains over 2K reasoning questions with different difficulty parameters.
+The dataset also includes the generation code, which can, in principle, be used to generate an unbounded number of reasoning questions.
+We evaluate our dataset on seven widely-used LLMs: GPT-5.4 (gpt5), o3 (o3), Llama3-70B  (llama) , Gemma3-27B  (gemma), DeepSeek-V3  (v3), GPT-4o  (gpt4) , and DeepSeek-R1-Distill-Llama-70B  (r1).
 For each LLM, we test whether it can produce valid and high-quality hypotheses.
-We also test whether popular reasoning-enhancing techniques, such as in-context learning and RLVR, which primarily were meant to improve deductive reasoning, can also benefit inductive and abductive reasoning.
+We also test whether popular reasoning-enhancing techniques, such as in-context learning and RLVR, which primarily were meant to improve deductive reasoning, can also improve inductive and abductive reasoning.
 Our findings can be summarized as follows:
@@
   2. 2.
-    The accuracy drops significantly when the ontology tree structure is complex and there are multiple hypotheses.
+    The accuracy drops significantly when the ontology tree structure is complex and contains multiple hypotheses.
   3. 3.
-    In some cases, LLMs may produce valid hypotheses (hypotheses which can explain all observations); but fail to produce high-quality hypotheses.
+    In most cases, LLMs may produce valid hypotheses (which can explain all observations); but fail to produce high-quality hypotheses (which adhere to Occam’s Razor).
   4. 4.
-    In general, we observe that reasoning-enhancing techniques for deductive reasoning only moderately benefits inductive and abductive reasoning.
+    In general, we observe that reasoning-enhancing techniques moderately improves inductive and abductive reasoning.
 2 Related Work
+Evaluating non-deductive reasoning capabilities.
+Dataset Controls for complexity of world models Natural language description Tests for inductive reasoning Test for abductive reasoning Evaluation of hypotheses quality Multiple hypotheses
+ART (anll) non-quantitative
+[…]
+AbductionRules (abductionrule)
+Wino-Why (winowhy)
+[…]
+UniADILR (Uni)
+LogiEval (logieval)
+MME-Reasoning (mme_reasoning)
+InAbHyD(ours)
+Table 1: Comparison of existing datasets to evaluate inductive and abductive reasoning capabilities.
+Table 1 summarizes the similarities and differences between our work and others on evaluating non-deductive reasoning capabilities of both LLMs and earlier models.
+Representative works on abductive reasoning evaluation, such as ART (anll), AbductionRules (abductionrule), and Wino-Why (winowhy), craft abductive reasoning questions from web sources, and ask LMs to explain everyday events.
+However, this approach confounds the problem of commonsense knowledge acquisition with the problem of abductive reasoning.
+In contrast, our framework provides all required facts in the context of each prompt, thereby evaluating LM’s abductive reasoning capabilities in a more controlled fashion.
+Representative works on inductive reasoning evaluation, such as InductiveBench (inductivebench) and SyGus (sygus) focus exclusively on highly-structured domains such as string translation and program synthesis.
+However, natural language-based inductive reasoning is also essential and is underexplored.
 Evaluating deductive reasoning capabilities.
-There is no consensus on whether LLMs can perform general deductive reasoning or are stochastic parrots (Bender et al., 2021).
-Researchers have designed different datasets and evaluation methodologies to answer this question  (Saparov et al., 2023; Saparov and He, 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024).
-LogiQA  (Liu et al., 2020) is a multiple-choice QA dataset to test multiple types of deductive reasoning of language models.
-However, it only attends to label accuracy (the correctness of the final answer) without considering the correctness of intermediate proof steps.
-In contrast, PrOntoQA  (Saparov and He, 2023) and PrOntoQA-OOD  (Saparov et al., 2023) evaluate both label accuracy and step-by-step proof correctness.
+There is no consensus on whether LLMs can perform general deductive reasoning or are stochastic parrots (parrot).
+Researchers have designed different datasets and evaluation methodologies to answer this question  (prontoood; prontoqa; logiqa; proofwriter; follo).
+LogiQA  (logiqa) is a multiple-choice QA dataset to test multiple types of deductive reasoning of language models.
+However, it focuses solely on label accuracy (the correctness of the final answer) without considering the correctness of intermediate proof steps.
+In contrast, PrOntoQA  (prontoqa; prontoood) evaluates both label accuracy and step-by-step proof correctness.
 They can synthetically generate proofs with different heights, widths, and deduction rules.
-Our work is similar to PrOntoQA and ProntoQA-OOD in spirit.
+Our work is similar to PrOntoQA in spirit.
 We both use a first-order ontology tree as the world model.
-However, PrOntoQA and PrOntoQA-OOD do not evaluate inductive and abductive reasoning.
-Evaluating non-deductive reasoning capabilities.
-[…]
-Dataset Controls for complexity of world models Natural language description Inductive reasoning Abductive reasoning Evaluation of hypotheses quality Multiple hypotheses
-ART (Zhao et al., 2023) non-quantitative
-[…]
-AbductionRules (Young et al., 2022)
-Wino-Why (Zhang et al., 2020)
-[…]
-[…]
-UniADILR (Sheng et al., 2025)
-Mirage (Li et al., 2025)
-Case2Code (Shao et al., 2024)
-InAbHyD (ours)
-Table 1 summarizes the similarities and differences between our work and others on evaluating non-deductive reasoning capabilities in both the pre- and LLM eras  (Young et al., 2022; Hua et al., 2025; Zhao et al., 2023; Yang et al., 2024; Zhang et al., 2021).
-Our work differs from others in the following ways.
-First, we focus on inductive and abductive reasoning and their interactions with deductive reasoning, while other works focus only on a single reasoning modality.
-Second, our reasoning problem is built upon a fictional world model with natural language descriptions.
-The fictional world model avoids training data contamination compared with ART  (Zhao et al., 2023), where the problem is directly taken from a web corpus; the natural language descriptions of the reasoning problem render the evaluation more practical compared with symbolic reasoning, such as in InductiveBench  (Hua et al., 2025).
-Third, our dataset is programmable with highly-controllable difficulty parameters, allowing us to evaluate LLMs’ reasoning capabilities under various scenarios.
+However, PrOntoQA does not evaluate non-deductive reasoning.
 Techniques to enhance reasoning.
-A large body of work has focused on improving the (deductive) reasoning capabilities of LLMs and related downstream tasks such as mathematics and coding  (Li, 2025; Wei et al., 2022; Yao et al., 2023; Wang et al., 2023; Su et al., 2025; Lambert et al., 2024; Zhou et al., 2022; Dong et al., 2024).
-The first line of work aims to elicit the emergent capabilities of LLMs by different formats of prompts, such as chain-of-thought (CoT)  (Wei et al., 2022), tree-of-thought (ToT)  (Yao et al., 2023), and self-consistency  (Wang et al., 2023).
-The second line of work leverages in-context learning by providing few-shot examples to LLMs.
-The third line of work views LLM token generation as a reinforcement learning problem and utilizes verifiable reward functions (reinforcement learning with verifiable reward, RLVRLambert et al.
-2024; Su et al.
-2025) to steer LLMs to improved reasoning.
-The fourth line of work scales LLMs’ test time to increase the accuracy of reasoning problems.
-It is worth noting that most techniques discussed above focus mainly on deductive reasoning.
+A large body of work has focused on improving the (deductive) reasoning capabilities of LLMs and related downstream tasks such as mathematics and coding  (ttt1; CoT; ToT; consistency; RLVRmultiple; tulu; ICL1; ICL2).
+The first line of work aims to elicit the emergent capabilities of LLMs using different prompt formats, such as chain-of-thought  (CoT; CoT), tree-of-thought  (ToT; ToT), self-consistency  (consistency), and in-context learning.
+The second line of work post-trains LLMs with reinforcement learning with verifiable reward (RLVR)  tulu; RLVRmultiple to improve LLM reasoning.
+Since non-deductive reasoning is inherently ambiguous, it is difficult to define a verifiable reward in non-deductive settings.
+The third line of work uses neuro-symbolic approaches to enhance the efficiency and reliability of LLM reasoning (ns1; ns2; ns3).
+However, they are either not directly applicable to unstructured natural language input or designed for deductive reasoning applications.
 This work investigates whether the above methods benefit inductive and abductive reasoning.
 3 Approach
-In this section, we describe how we generate the dataset and evaluate LLMs on this dataset.
-At a high level, the dataset consists of synthetically-generated reasoning examples in natural language.
-The question description is a user prompt, and the format specification is a system prompt to LLMs.
-We parse LLMs’ output to evaluate their inductive and abductive reasoning capabilities on metrics such as accuracy.
-We acknowledge that the potential space of inductive and abductive reasoning examples is vast, so we restrict ourselves to what is expressible in first-order logic.
-By composing the inductive and abductive reasoning steps in Table 2, our evaluation framework can readily cover all inductive and abductive reasoning questions in first-order logic.
+3.1 Overview
+In this section, we describe how we generate the dataset and evaluate LLMs.
+The dataset consists of synthetically-generated reasoning examples in natural language.
+Figure 2 illustrates the pipeline for generating each question.
+We evaluate LLM’s inductive and abductive reasoning capabilities on metrics such as accuracy and their adherence to Occam’s Razor.
+By composing inductive and abductive reasoning steps shown in Table 2, our evaluation framework can readily be extended to produce any abductive/inductive reasoning question expressible in first-order logic.
+Step 1 (§3.2) Generate the world model Step 2 (§3.3) Generate observations Step 3: Naturalistic Question Generation (§3.4) Convert theories and observations into natural language using templates Use GPT-4o to paraphrase examples into more naturalistic language
+Figure 2: Pipeline for generating each reasoning question in InAbHyD.
 Each reasoning example contains a fictional world model and a set of observations.
@@
   2. 2.
-    Second, we traverse the ontology tree and randomly mark some axioms as hidden. Each hidden axiom is the ground truth for hypotheses.
+    Second, we traverse the ontology tree and randomly mark some axioms as hidden. Each hidden axiom is the ground truth.
   3. 3.
-    Third, for each hidden axiom, we generate observations so that LLMs have enough information to produce ground truth hypotheses.
+    Third, for each hidden axiom, we generate observations so that LLMs have enough information to produce ground-truth hypotheses.
   4. 4.
-    Fourth, we translate the incomplete world model and observations into natural language.
-Subtask Reasoning type tested World model Observations Ground truth hypothesis Proof trees for observations
+    Fourth, we transform the incomplete world model and observations into natural language.
+Subtask Reasoning type tested World model Observations Ground-truth hypothesis Proof trees for observations
 Infer property Inductive o(A)
@@
 […]
-Table 2: An overview of various subtasks in InAbHyD and their details in the simplest case (a height-1 ontology tree with a single hypothesis).
+Table 2: An overview of various subtasks and their details in the simplest case (a height-1 ontology tree with a single hypothesis).
 Here o denotes an ontology concept (e.g, cat) and o(A) denotes the membership (e.g., cat(Sally) means Sally is a cat).
@@
 We only show the proof tree for one observation for simplicity; other observations can be derived similarly.
-We show hypotheses green and axioms in world models purple in the proof tree.
+In the proof tree, we show hypotheses in green and axioms in world models in purple.
 See Table 2 for an overview of the different subtasks and their details in the simplest case (a height-1 ontology tree with a single hypothesis).
@@
   1. 1.
-    Strong accuracy: Hypotheses are strongly correct if they exactly match ground truth hypotheses.
+    Strong accuracy: Hypotheses are strongly correct if they exactly match ground-truth hypotheses (simplest hypothesis).
   2. 2.
@@
   3. 3.
-    Hypothesis quality: A quantitative measurement of the quality of the LLMs’ produced hypotheses based on the principles of parsimony.
+    Hypothesis quality: A quantitative measurement of the quality of the LLMs’ produced hypotheses based on Occam’s Razor.
 Next, we describe each step in more detail.
-World model generation
+3.2 World model generation
 In our approach, the world model is an ontology tree in first-order logic.
@@
 cat) with properties (e.g.
-All cats are cute, or equivalently in logical form $\forall x(\texttt{cat}(x)\rightarrow\texttt{cute}(x))$) and members (e.g.
-Amy is a cat, or equivalently in logical forms, cat(Amy)).
-Each edge in the ontology tree denotes a subtype relationship (e.g., the ragdoll node being a child of the cat node implies Each ragdoll is a cat, or equivalently in logical form, $\forall x(\texttt{ragdoll}(x)\rightarrow\texttt{cat}(x))$).
+All cats are cute ) and members (e.g.
+Amy is a cat).
+Each edge in the ontology tree denotes a subtype relationship (e.g., the ragdoll node being a child of the cat node implies Each ragdoll is a cat).
 We first generate the topology of the ontology tree with the height as a controllable parameter.
-Each non-leaf node has two or three child nodes.
+Each non-leaf node has a set of child nodes.
 The height of the ontology tree is a variable that characterizes the complexity of the world model.
-We then make two passes of the tree in level order traversal.
-In the first pass, for each node, we allocate a concept name, three properties, [^4] The properties can also be negated properties, e.g., All cats are not slow.
-and three members.
+We then make two passes over the tree using level-order traversal.
+In the first pass, for each node, we allocate a concept name, properties, and members  [^3] The properties can also be negated, e.g., All cats are not slow, which corresponds to the logical form $\forall x(\texttt{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}cat}}(x)\rightarrow\neg\texttt{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}slow}}(x))$ (negation always takes narrow scope).
+The number of properties, members, and child nodes follows a zipf distribution with $a=3$..
 After the first pass, we have a complete world model with three types of axioms: (1) properties (e.g., All cats are cute), (2) membership relations (e.g., Amy is a cat), (3) subtype relations (e.g., Each ragdoll is a cat).
 In the second pass, we randomly mark some axioms as hidden.
-These axioms are the ground truth hypotheses.
+These axioms are the ground-truth hypotheses.
 After the second pass, we have an incomplete world model.
-Observation generation
-Next, we generate observations so that LLMs have sufficient information to produce the ground truth hypotheses.
-The observations must be complex enough that the reasoning example is non-trivial, yet they should also ensure the reasoning is feasible.
-  • To generate an example where the task is to infer a property (e.g., All cats are cute), we first identify the type in the ontology with this property (e.g., cat), then we sample members of this type and subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where the members have this property (e.g., Amy is cute; Sally is cute; Polly is cute).
-  • To generate an example where the task is to infer a membership relation (e.g., Amy is a cat), we first identify the type in the ontology with this member (e.g., cat), then we sample properties of this type and supertypes (e.g., All cats are cute; All felines are not slow; All mammals are hairy), and add observations where this member have these properties (e.g., Amy is cute; Amy is not slow; Amy is hairy).
-  • To generate an example where the task is to infer a subtype relation (e.g., All cats are felines), we first identify the subtype in the ontology (e.g., cat), then we sample members of this type and its subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where these members belong to the supertype (e.g., Amy is a feline; Polly is a feline; Sally is a feline).
-Translate theories and observations into natural language
-We use the same grammar as in PrOntoQA for translating logical forms into natural language.
-Properties are translated into the form: All cats are cute; subtype relations are randomly translated into one of three forms: Each ragdoll is a cat; Every ragdoll is a cat; All ragdolls are cats; membership relations are translated into the form Amy is a cat.
-Hypothesis quality
-We define a metric to evaluate the quality of hypotheses based on the principles of parsimony, or Occam’s Razor (i.e., the priciple that, other things being equal, simpler theories are better;  Baker 2022).
+3.3 Observation generation
+Next, we generate observations so that LLMs have sufficient information to produce the ground-truth hypotheses.
+The observations must be complex enough that the reasoning example is non-trivial, yet they should also ensure the problem is feasible.
+  • To generate an example where the task is to infer a property (e.g., All cats are cute), we first identify the type in the ontology with this property (e.g., cat), then we sample members of this type and subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where the members have this property (e.g., Amy is cute; Sally is cute;).
+  • To generate an example where the task is to infer a membership relation (e.g., Amy is a cat), we first identify the type in the ontology with this member (e.g., cat), then we sample properties of this type and supertypes (e.g., All cats are cute; All felines are not slow; All mammals are hairy), and add observations where this member have these properties (e.g., Amy is cute; Amy is not slow).
+  • To generate an example where the task is to infer a subtype relation (e.g., All cats are felines), we first identify the subtype in the ontology (e.g., cat), then we sample members of this type and its subtypes (e.g., Amy is a cat; Sally is a ragdoll; Polly is a shorthair), and add observations where these members belong to the supertype (e.g., Amy is a feline; Polly is a feline).
+3.4 Naturalistic Question Generation
+We first use the same grammar as in PrOntoQA for converting logical forms into natural language using templates.
+Properties are converted into the form: All cats are cute; subtype relations are randomly converted into one of three forms: Each ragdoll is a cat; Every ragdoll is a cat; All ragdolls are cats; membership relations are converted into the form Amy is a cat.
+We then use GPT-4o to paraphrase it into more natural language.
+Details are provided in Appendix D.
+3.5 Hypothesis quality
+We define a metric to evaluate the quality of hypotheses based on the principles of Occam’s Razor (i.e., the priciple that, other things being equal, simpler theories are better;  occ).
 In this regard, we encourage one hypothesis to explain as many observations as possible and penalize unnecessary hypotheses.
-Formally, given a world model $\mathcal{T}$, a set of observations $\mathcal{O}$, their associated proof trees $\mathcal{P}$,candidate hypotheses $\mathcal{H}$, and ground truth hypotheses $\mathcal{H}^{*}$, for $h\in\mathcal{H}$, $n(h)$ denotes the number of times $h$ appears in the proof trees of $\mathcal{O}$ (i.e., how many times $h$ is used as a premise of a proof step).
-We use $|\mathcal{H}|$ to denote the cardinality of $\mathcal{H}$.
+Formally, given a world model $\mathcal{T}$, a set of observations $\mathcal{O}$, their associated proof trees $\mathcal{P}$,candidate hypotheses $\mathcal{H}$, and ground-truth hypotheses $\mathcal{H}^{*}$, for $h\in\mathcal{H}$, $n(h)$ denotes the number of times $h$ appears in the proof trees of $\mathcal{O}$ (i.e., how many times $h$ is used as a premise of a proof step).
+We first use $|\mathcal{H}|$ to denote the cardinality of $\mathcal{H}$.
 The quality of $\mathcal{H}$, $q(\mathcal{H})$ is defined by  [^4] We write the equation as the ratio of two averages: the numerator is the average accuracy of the predicted hypothesis (i.e., the average number of times a hypothesis appears in any valid proof), and the denominator is the average number of times a hypothesis appears in the shortest (i.e., most parsimonious) proof.:
 […]
-According to the definition, the ground truth hypotheses $\mathcal{H}^{*}$ have a quality score of 1.
+According to the definition, the ground-truth hypotheses $\mathcal{H}^{*}$ have a quality score of 1.
 Consider the example in Figure 1, here $\mathcal{H}^{*}$ is {Fae is a tiger, All mammals are hairy, All rodents are mammals} and $|\mathcal{H}^{*}|=3$.
@@
 In this case, $q(\mathcal{H})=\frac{3\times 10}{5\times 9}=0.67$.
-Appendix C contains the complete algorithms for generating reasoning examples and evaluating hypotheses.
-4 Results
+We further motivate the hypothesis metric using Bayesian decision theory in Appendix E.1.
+We compare the metric to human evaluation of hypotheses in Appendix E.2.
+Appendix F contains the complete algorithms for generating reasoning examples and evaluating hypotheses.
+4 Experiments
 4.1 Experimental setup
@@
 We evaluate hypotheses on three metrics: weak accuracy, strong accuracy, and quality.
-We compute 95% confidence intervals for both accuracy metrics using the Wilson interval (Wilson, 1927).
-We use the OpenAI API to evaluate GPT-4o and together.ai platform to evaluate Llama3-70B, Gemma3-27B, DeepSeek-V3, and DeepSeek-R1-Distill-Llama-70B.
-In our experiments, we vary the following variables: (1) Different types of tasks (i.e., infer property, infer membership relation, or infer subtype relation), (2) Single hypothesis or multiple hypotheses, (3) The height of the ontology tree, (4) Whether the reasoning example contains in-context demonstrations and whether they are in-distribution or out-of-distribution.
+We compute 95% confidence intervals for both accuracy metrics using the Wilson interval (wilson).
+We use the OpenAI API and together.ai platform.
+In our experiments, we vary the following variables: (1) Different types of tasks (i.e., infer property, infer membership relation, or infer subtype relation), (2) Single hypothesis or multiple hypotheses, (3) The height of the ontology tree, (4) Whether the reasoning example contains in-context demonstrations and whether they are in-distribution (5) Different reasoning-augmented prompts such as buffer-of-thought (bot).
 4.2 Zero-shot results
-To begin evaluating the inductive and abductive reasoning abilities of LLMs, we first investigate the simplest setting with zero-shot prompting.
+We first investigate the simplest setting with zero-shot prompting.
 4.2.1 Single hypothesis
@@
 For example, even under a height-4 ontology tree, Gemma3-27B can achieve near 50% weak accuracy, while the number drops to around 10% for infer subtype relation tasks.
-This may be because the weak correctness for infer membership relation tasks only requires searching for the concept that has some property, but do not need to understand the structure of the ontology tree.666e.g., given Amy is cute, LLMs only need to search for a concept which has the property cute to infer the membership relation of Amy.
+This may be because the weak correctness for infer membership relation tasks only requires searching for the concept that has some property, but do not need to understand the structure of the ontology tree.555e.g., given Amy is cute, LLMs only need to search for a concept which has the property cute to infer the membership relation of Amy.
 However, this is not the case for infer subtype relation and infer properties tasks.
-We observe a similar trend for infer subtype relation and infer properties tasks, as we can view the concept’s parent as a special property of that concept.
-1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00Quality1234Height0.000.250.500.751.00Weak Accuracy1234Height0.000.250.500.751.00Strong Accuracy1234Height0.000.250.500.751.00QualityTask: infer membership relationTask: infer propertyTask: infer subtype relationGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
+We observe a similar trend for infer subtype relation and infer properties tasks, as we can view the concept’s parent as a special property.
+12340.00.20.40.60.81.0Weak Accuracy12340.00.20.40.60.81.0Strong Accuracy12340.00.20.40.60.81.0Quality12340.00.20.40.60.81.0Weak Accuracy12340.00.20.40.60.81.0Strong Accuracy12340.00.20.40.60.81.0Quality1234Height0.00.20.40.60.81.0Weak Accuracy1234Height0.00.20.40.60.81.0Strong Accuracy1234Height0.00.20.40.60.81.0QualityTask: infer membership relationTask: infer propertyTask: infer subtype relationGPT-4oLlama3-70BDeepSeek-V3Gemma3-27B
 Figure 3: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with single hypothesis.
@@
 Each reasoning example can contain multiple hypotheses.
-To check the difficulty of each reasoning example, we sample the ground truth hypotheses such that the number of hypotheses grows linearly w.r.t.
+To check the difficulty of each reasoning example, we sample the ground-truth hypotheses such that the number of hypotheses grows linearly w.r.t.
 the height of the ontology tree.
-[^7] Each ground truth hypothesis has $\sim 3$ observations.
+[^6] Each ground-truth hypothesis has $\sim 3$ observations.
 So the number of observations also grows linearly w.r.t.
@@
 the height of the ontology tree.
-See Table 3 for statistics under multi-hypothesis scenarios and the top row of Figure 3 for results.
-The height of the ontology tree # of world model axioms # of observations # of ground truth hypotheses 1 9.0 10.0 3.0 2 14.0 11.8 3.5 3 25.5 15.2 4.6 4 46.8 20.0 6.6
-Table 3: Basic statistics of multi-hypothesis examples in InAbHyD .
-We list the average number of world model axioms, observations, and ground truth hypotheses w.r.t.
-the height of the ontology tree in the same set of reasoning examples we used in evaluation.
-Note that the number of the ground truth hypotheses and the number of observations grow linearly with the height of the ontology tree.
-In contrast, the number of world model axioms grows exponentially.
-In general, the performance in the multi-hypothesis setting differed significantly from that in the single-hypothesis setting.
+See Table 3 in Appendix B for statistics under multi-hypothesis scenarios and the top row of Figure 4 for results.
+Performance in the multi-hypothesis setting differed significantly from that in the single-hypothesis setting.
 For example, the accuracy of all models drops significantly when the height of the ontology tree is increased from 1 to 2, with accuracy going from above 80% to below 50%, except for GPT-4o.
-This is surprising because the average number of ground truth hypotheses only increases from $3.0$ to $3.5$.
+This is surprising because the average number of ground-truth hypotheses increases only from $3.0$ to $3.5$.
 We attribute this to the coupling effect of multiple hypotheses to produce and the topology of the ontology tree.
-Reasoning examples under the ontology tree with height 3 or 4 achieve lower accuracies and quality scores for all models.
+Reasoning examples under the ontology tree with heights 3 or 4 achieve lower accuracy and quality scores across all models.
 We also observe a notable gap between weak accuracy and quality.
-For example, under the height-4 ontology tree, all models can achieve at least 20% weak accuracy, but struggle with extremely low quality and strong accuracy scores.
+For example, under the height-4 ontology tree, all models can achieve at least 20% weak accuracy but struggle with extremely low-quality and strong-accuracy scores.
 We manually inspect some incorrect hypotheses and summarize the most common types of errors in Section 4.5.
@@
 […]
-Figure 3: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-short (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom).
-We list the absolute results of the 8-shot in-distribution and out-of-distribution in Appendix D.
-LLMs have shown extraordinary generalization capabilities with few-shot examples coupled with CoT prompts (Brown et al., 2020).
-We are also interested in whether in-context learning can help with inductive and abductive reasoning.
-We add eight demonstrations to each reasoning example, with the demonstrations’ ground truth hypotheses and CoTs.
+Figure 4: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-shot (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom).
+We list the absolute results of the 8-shot in-distribution and out-of-distribution in Appendix G.
+LLMs have shown extraordinary generalization capabilities with few-shot examples coupled with CoT prompts (iclclassicla).
+We are also interested in whether in-context learning can help with inductive/abductive reasoning.
+We add $8$ demonstrations to each example, with the demonstrations’ ground-truth hypotheses and CoTs.
 The CoT consists of each observation’s step-by-step proof.
@@
 In both settings, we ensure demonstrations cover all types of tasks.
+We present our results in Figure 4.
 The in-distribution demonstrations are slightly more helpful than out-of-distribution demonstrations.
-This result is consistent with the hypothesis that in-context learning mimics Bayesian inference (Xie et al., 2022).
+This result is consistent with the hypothesis that in-context learning mimics Bayesian inference (cotbay).
 We also observe that in-context learning with in-distribution demonstrations improves strong accuracy and quality in ontology trees with heights 3 and 4.
@@
 In contrast, we do not observe a significant performance improvement from out-of-distribution demonstrations.
-4.4 Does RLVR help?
-1234Height0.00.20.40.60.81.0Weak Accuracy1234Height0.00.20.40.60.81.0Strong Accuracy1234Height0.00.20.40.60.81.0QualityLlama3-70BDeepSeek-R1-Distill-Llama-70B
-Figure 4: Weak accuracy, strong accuracy, and quality versus the height of the ontology tree comparison results for Llama3-70B and DeepSeek-R1-Distill-Llama-70B with multiple hypotheses.
-RLVR has demonstrated significant improvements in other reasoning benchmarks, and so we test whether this improvement generalizes to abductive and inductive reasoning.
-Specifically, we compare Llama3-70B and DeepSeek-R1-Distill-Llama-70B.
-In general, we observe an average performance improvement between 10% and 20%.
-To analyze why DeepSeek-R1-Distill-Llama-70B performs better than Llama3-70B, we also manually inspect the reasoning output of DeepSeek-R1-Distill-Llama-70B (i.e., tokens that are inside <think> and </think>).
-After producing hypotheses, DeepSeek-R1-Distill-Llama-70B will usually attempt to verify those hypotheses.
-This verification process is essentially a deductive reasoning process, and can catch some reasoning errors or fix low-quality hypotheses.
-A complete example is in Appendix E.
+4.4 Do reasoning models and reasoning-augmented prompts help?
+Figure 5: Results for Llama3-70B and two large reasoning models gpt-5.4 and o3 with chain-of-thought prompting (top) and buffer-of-thought prompting (bottom).
+Recent developments in large reasoning models, particularly those trained with RLVR, have led to substantial improvements across many evaluations.
+As shown in Figure 5, stronger models such as gpt5.4 and o3 consistently outperform LLaMA3-70B in both weak and strong accuracy and overall quality across different ontology tree heights.
+Moreover, reasoning-oriented prompting strategies further enhance performance: buffer-of-thought (BoT; bot) prompting generally yields higher scores than chain-of-thought prompting (CoT), especially in weak accuracy.
+However, despite these gains, a persistent gap remains between weak and strong accuracy across all models and settings—–for instance, weak accuracy often approaches near-perfect levels at lower heights, while strong accuracy remains notably lower and degrades more rapidly as height increases.
+This suggests that although recent advances and prompting techniques significantly improve reasoning capabilities, high-quality hypothesis generation remains an open challenge.
 4.5 Analysis of common errors
@@
 We summarize some of the most common types of errors here for further insights.
-We provide example responses in each category in Appendix E.
+We provide example responses in each category, quantitative analysis, and the details of the manual inspection process in Appendix H.
 (1) Wrong ontology direction.LLMs sometimes produce hypotheses in the wrong direction.
-For example, the ground truth hypothesis is All cats are mammals, while LLMs will produce the hypothesis All mammals are cats.
+For example, the ground-truth hypothesis is All cats are mammals, while LLMs will produce the hypothesis All mammals are cats.
 (2) Ignore the ontology and produce unnecessary hypotheses.LLMs sometimes can produce unnecessary hypotheses because they ignore the ontology.
-For example, the ground truth hypothesis is All cats are cute, while LLMs produce two hypotheses: All cats are cute; All ragdolls are cute.
 (3) Fall back to trivial hypotheses.Technically, the observation is also a valid hypothesis, but it is considered low-quality because it can only explain itself.
@@
 Usually, this happens when the entity appears in the in-context demonstrations, but LLMs use them in the test question.
-(5) Confuse members with concepts.We use fictional names for concepts and real-person names for members.
-However, LLMs sometimes do confuse these two.
-For example, the ground truth hypothesis is Amy is a wumpus, while LLMs produce Wumpus is Amy.
-5 Conclusions and future work
-In this study, we introduced a programmable and synthetic dataset, InAbHyD .
-Our dataset serves as an out-of-distribution evaluation of the inductive and abductive reasoning abilities of LLMs.
-We do not claim that success in our benchmark automatically implies success in real-world tasks, but rather, our benchmark serves as a checkpoint for developing models that are capable of robust real-world abductive and inductive reasoning.
-We systematically studied LLMs’ inductive and abductive reasoning capabilities.
-Our results suggest that current LLMs can’t generate high-quality hypotheses that match human-level intelligence.
-For example, suppose we use LLMs as a scientific discovery assistant.
-In that case, they may generate superficial hypotheses that can explain observations but fail to uncover more general hypotheses that reveal hidden structures or connections in science.
-Our results can serve to better inform practitioners of LLMs of their limitations, especially in applications that rely heavily on inductive and abductive reasoning.
-The dataset itself opens new avenues for future work, such as training or fine-tuning models to improve their inductive and abductive reasoning capabilities.
-One limitation of this work is that first-order logic (FOL) limits the expressiveness of the world model and the reasoning example.
-We leave it as a future work to use higher-order logic (HOL) to generate more complex world models and reasoning examples, which would require translating from HOL into natural language and vice versa.
-The other crucial future direction is improving the quality of LLMs’ hypotheses.
-For that, LLMs may benefit from RLVR  (Lambert et al., 2024) that explicitly conveys principles of parsimony in inductive and abductive reasoning with an appropriate reward.
-Another interesting future direction will be to determine whether large language models fine-tuned on this synthetic dataset (using supervised fine-tuning or reinforcement learning from verifiable rewards) could improve their performance on real-world tasks involving inductive and abductive reasoning.
-If this is the case, this synthetic dataset is more valuable because we can essentially generate an infinite number of synthetic data points, without worrying about the bottleneck of sampling data from the real world.
-Appendix A Detailed explanations of Figure 1
-Here we list proof trees for each observation in Figure 1. Hypotheses are in green and world model axioms are in purple.
-Fae is strong
-[…]
-Fae is not slow
-$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\neg\texttt{slow}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\neg\texttt{slow}(x))$     $\texttt{tiger}(\texttt{Fae})$                             $\neg\texttt{slow}(\texttt{Fae})$
-Fae is warm-blooded
-$\forall x(\texttt{tiger}(x)\rightarrow\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\rightarrow\texttt{mammal}(x))$           $\forall x(\texttt{tiger}(x)\rightarrow\texttt{mammal}(x))$     $\forall x(\texttt{mammal}(x)\rightarrow\texttt{warm-blooded}(x))$                              $\forall x(\texttt{tiger}(x)\to\texttt{warm-blooded}(x))$     tiger(Fae)                                                     warm-blooded(Fae)
-Sam is hairy
-$\forall x(\texttt{mammal}(x)\to\texttt{hairy}(x))$     mammal(Sam)           hairy(Sam)
-Alice is hairy
-$\forall x(\texttt{cat}(x)\to\texttt{feline}(x))$     $\forall x(\texttt{feline}(x)\to\texttt{mammal}(x))$           $\forall x(\texttt{cat}(x)\to\texttt{mammal}(x))$     cat(Alice)                           mammal(Alice)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Alice)
-Bob is hairy
-$\forall x(\texttt{dog}(x)\to\texttt{canidae}(x))$     $\forall x(\texttt{canidae}\to\texttt{mammal}(x))$            $\forall x(\texttt{dog}(x)\to\texttt{mammal}(x))$     dog(Bob)                             mammal(Bob)     $\forall x(\texttt{mammal}(x)\to\texttt{hariy}(x))$                                                hairy(Bob)
-Jack is a mammal
-rat(Jack)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Jack)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Jack)
-Noah is a mammal
-rat(Noah)     $\forall x(\texttt{rat}(x)\to\texttt{rodent}(x))$           rodent(Noah)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                            mammal(Noah)
-Oliver is a mammal
-squirrel(Oliver)     $\forall x(\texttt{squirrel}(x)\to\texttt{rodent}(x))$             rodent(Oliver)      $\forall x(\texttt{rodent}(x)\to\texttt{mammal}(x))$                                  mammal(Oliver)
-Appendix B More examples
-Single hypothesis; task: infer property; height-1 ontology tree.
-World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
-Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
+5 Conclusion
+In this study, we systematically examined the inductive/abductive reasoning capabilities of LLMs, especially whether they follow Occam’s Razor.
+Our results suggest that current LLMs can’t generate high-quality hypotheses that match human-level parsimony.
+Our results can serve to better inform practitioners of LLMs of their limitations, especially in applications that rely heavily on inductive/abductive reasoning.
+The proposed dataset opens new avenues for future work, such as training models to improve their reasoning capabilities.
+Limitations
+Our synthetic dataset serves as an out-of-distribution evaluation of the inductive and abductive reasoning abilities of LLMs.
+We do not claim that success in our benchmark automatically implies success in real-world tasks.
+Still, our benchmark serves as a checkpoint for developing models that are capable of robust real-world abductive and inductive reasoning.
+Another limitation of this work is that first-order logic (FOL) restricts the expressiveness of the world model and the reasoning example.
+We leave it as future work to utilize higher-order logic (HOL) to generate more complex world models and reasoning examples.
+Appendix A Preliminary background
+A.1 Examples of different types of reasoning
+For example, given “All humans are moral” and “Socrates is a human”, we can conclude that “Socrates is moral”.
+For example, if we see many cute cats, we may inductively reason that all cats are cute.
+For example, if we hear a bird sound outside, we may abductively reason that there is a bird.
+The latter two types of reasoning share a common high-level goal: to find explanations for observations.
+Furthermore, the composition of inductive and abductive reasoning is prevalent in real-world scenarios.
+For example, given (i) Amy has a skin rash and leg pain and has been diagnosed with a new, unknown virus.
+(ii) Sam has a skin rash and leg pain and has been diagnosed with the same new unknown virus.
+(iii) Tom also has a skin rash and leg pain.
+We can first induce the rule: The new unknown virus causes skin rash and leg pain, and then we can abduce that Tom is infected with this new virus.
+A.2 Proof trees
+A proof tree is a tree-structured representation of a logical derivation where each node corresponds to a statement, and each edge represents the application of an inference rule.
+The root of the tree is the final conclusion, while the leaves correspond to premises or axioms.
+Proof trees are commonly used in logic, theorem proving, and formal reasoning to make the structure of derivations explicit.
+Formally, let $\Gamma$ denote a set of premises and $\varphi$ denote a conclusion.
+A proof tree demonstrates that $\Gamma\vdash\varphi$ by recursively applying inference rules until all leaf nodes correspond to elements of $\Gamma$ or axioms.
+Example.
+Consider the following premises:
+$1.\quad$ $P\rightarrow Q$
+$2.\quad$ $Q\rightarrow R$
+$3.\quad$ $P$
+We can derive $R$ using two applications of modus ponens.
+We first derive $Q$ given the premises $P$ and $P\rightarrow Q$.
+Next, we derive $R$ using the premises $Q$ and $Q\rightarrow R$.
+This proof can be represented as a proof tree, which is shown below.
+$\frac{\frac{P\rightarrow Q\qquad P}{Q}\qquad Q\rightarrow R}{R}$
+The tree therefore makes explicit the hierarchical structure of the reasoning process.
+Appendix B Basic statistics of InAbHyD
+We list basic statistics of InAbHyD in Table 1.
+Ontology tree height # of world model axioms # of observations # of ground-truth hypotheses # of examples 1 9.0 10.0 3.0 500 2 14.0 11.8 3.5 500 3 25.5 15.2 4.6 500 4 46.8 20.0 6.6 500
+Table 3: Basic statistics of multi-hypothesis examples in InAbHyD.
+We list the average number of world model axioms, observations, and ground-truth hypotheses w.r.t.
+the height of the ontology tree in the same set of reasoning examples we used in evaluation.
+Note that the number of the ground-truth hypotheses and the number of observations grow linearly with the height of the ontology tree.
+In contrast, the number of world model axioms grows exponentially.
+Appendix C More examples of reasoning questions
+In this section, we provide additional reasoning examples generated using our proposed framework.
+For each setting listed below, we provide an example consisting of a world model, list of observations, and ground-truth hypothesis/hypotheses.
+Single hypothesis example; task: infer property; height-1 ontology tree.
+World Model: Amy is a dalpist.
+Pamela is a dalpist.
+Jerry is a dalpist.
+Observations: Amy is rainy.
+Jerry is rainy.
+Pamela is rainy.
 Hypotheses: Dalpists are rainy.
-Single hypothesis; task: infer property; height-2 ontology tree.
-World Model: Rompuses are dalpists. Sarpers are dalpists. Each gergit is a dalpist. Pamela is a sarper. Jerry is a dalpist. Amy is a rompus.
-Observations: Amy is rainy. Pamela is rainy. Jerry is rainy.
+Single hypothesis example; task: infer property; height-2 ontology tree.
+World Model: Rompuses are dalpists.
+Sarpers are dalpists.
+Each gergit is a dalpist.
+Pamela is a sarper.
+Jerry is a dalpist.
+Amy is a rompus.
+Observations: Amy is rainy.
+Pamela is rainy.
+Jerry is rainy.
 Hypotheses: Dalpists are rainy.
-Single hypothesis; task: infer property; height-3 ontology tree.
+Single hypothesis example; task: infer property; height-3 ontology tree.
 World Model: Pamela is a porpor.
@@
 Rompuses are dalpists.
-Observations: Jerry is not muffled. Amy is not muffled. Pamela is not muffled.
+Observations: Jerry is not muffled.
+Amy is not muffled.
+Pamela is not muffled.
 Hypotheses: Each dalpist is not muffled.
-Single hypothesis; task: infer property; height-4 ontology tree.
+Single hypothesis example; task: infer property; height-4 ontology tree.
 World Model: Amy is a felper.
@@
 Felpers are storpists.
-Observations: Amy is translucent. Pamela is translucent. Jerry is translucent.
+Observations: Amy is translucent.
+Pamela is translucent.
+Jerry is translucent.
 Hypotheses: Each dalpist is translucent.
-Single hypothesis; task: infer membership relation; height-1 ontology tree.
-World Model: Dalpists are liquid. Dalpists are rainy. Each dalpist is brown.
-Observations: Jerry is brown. Jerry is rainy. Jerry is liquid.
+Single hypothesis example; task: infer membership relation; height-1 ontology tree.
+World Model: Dalpists are liquid.
+Dalpists are rainy.
+Each dalpist is brown.
+Observations: Jerry is brown.
+Jerry is rainy.
+Jerry is liquid.
 Hypotheses: Jerry is a dalpist.
-Single hypothesis; task: infer membership relation; height-2 ontology tree.
+Single hypothesis example; task: infer membership relation; height-2 ontology tree.
 World Model: Rompuses are dalpists.
@@
 Dalpists are liquid.
-Observations: Pamela is rainy. Pamela is liquid. Pamela is brown.
+Observations: Pamela is rainy.
+Pamela is liquid.
+Pamela is brown.
 Hypotheses: Pamela is a gergit.
-Single hypothesis; task: infer membership relation; height-3 ontology tree.
+Single hypothesis example; task: infer membership relation; height-3 ontology tree.
 World Model: Each sarper is a dalpist.
@@
 Twimpees are not muffled.
-Observations: Nicole is moderate. Nicole is large. Nicole is not muffled.
+Observations: Nicole is moderate.
+Nicole is large.
+Nicole is not muffled.
 Hypotheses: Nicole is a twimpee.
-Single hypothesis; task: infer membership relation; height-4 ontology tree.
+Single hypothesis example; task: infer membership relation; height-4 ontology tree.
 World Model: Amy is a rompus.
@@
 Debra is a storpist.
-Observations: Michael is fruity. Michael is translucent. Michael is moderate.
+Observations: Michael is fruity.
+Michael is translucent.
+Michael is moderate.
 Hypotheses: Michael is a shumple.
-Single hypothesis; task: infer subtype relation; height-1 ontology tree.
-World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
-Observations: Amy is a rompus. Jerry is a rompus. Pamela is a rompus.
+Single hypothesis example; task: infer subtype relation; height-1 ontology tree.
+World Model: Amy is a dalpist.
+Pamela is a dalpist.
+Jerry is a dalpist.
+Observations: Amy is a rompus.
+Jerry is a rompus.
+Pamela is a rompus.
 Hypotheses: Dalpists are rompuses.
-Single hypothesis; task: infer subtype relation; height-2 ontology tree.
-World Model: Gergits are dalpists. Porpors are dalpists. Each sarper is a dalpist. Pamela is a porpor. Jerry is a dalpist. Amy is a gergit.
-Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
+Single hypothesis example; task: infer subtype relation; height-2 ontology tree.
+World Model: Gergits are dalpists.
+Porpors are dalpists.
+Each sarper is a dalpist.
+Pamela is a porpor.
+Jerry is a dalpist.
+Amy is a gergit.
+Observations: Amy is a rompus.
+Pamela is a rompus.
+Jerry is a rompus.
 Hypotheses: Dalpists are rompuses.
-Single hypothesis; task: infer subtype relation; height-3 ontology tree.
+Single hypothesis example; task: infer subtype relation; height-3 ontology tree.
 World Model: Amy is a storpist.
@@
 Gergits are dalpists.
-Observations: Jerry is a rompus. Amy is a rompus. Pamela is a rompus.
+Observations: Jerry is a rompus.
+Amy is a rompus.
+Pamela is a rompus.
 Hypotheses: Each dalpist is a rompus.
@@
 Porpors are dalpists.
-Observations: Amy is a rompus. Pamela is a rompus. Jerry is a rompus.
+Observations: Amy is a rompus.
+Pamela is a rompus.
+Jerry is a rompus.
 Hypotheses: Each dalpist is a rompus.
-Multiple hypothesis; height-1 ontology tree.
-World Model: Debra is a dalpist. Amy is a dalpist. Pamela is a dalpist. Edward is a dalpist. Dalpists are liquid. Each dalpist is moderate. Each dalpist is brown. Barbara is a dalpist.
-Observations: Barbara is rainy. Jerry is a rompus. Jerry is liquid. Jerry is brown. Pamela is a rompus. Edward is rainy. Debra is rainy. Jerry is moderate. Amy is a rompus.
-Hypotheses: Dalpists are rompuses. Dalpists are rainy. Jerry is a dalpist.
-Multiple hypothesis; height-2 ontology tree.
+Multiple hypothesis example; height-1 ontology tree.
+World Model: Debra is a dalpist.
+Amy is a dalpist.
+Pamela is a dalpist.
+Edward is a dalpist.
+Dalpists are liquid.
+Each dalpist is moderate.
+Each dalpist is brown.
+Barbara is a dalpist.
+Observations: Barbara is rainy.
+Jerry is a rompus.
+Jerry is liquid.
+Jerry is brown.
+Pamela is a rompus.
+Edward is rainy.
+Debra is rainy.
+Jerry is moderate.
+Amy is a rompus.
+Hypotheses: Dalpists are rompuses.
+Dalpists are rainy.
+Jerry is a dalpist.
+Multiple hypothesis example; height-2 ontology tree.
 World Model: Nicole is a gergit.
@@
 Amy is a rompus.
-Hypotheses: Dalpists are rompuses. Dalpists are liquid. Every gergit is moderate. Amy is a porpor.
-Multiple hypothesis; height-3 ontology tree.
+Hypotheses: Dalpists are rompuses.
+Dalpists are liquid.
+Every gergit is moderate.
+Amy is a porpor.
+Multiple hypothesis example; height-3 ontology tree.
 World Model: Barbara is a sarper.
@@
 Pamela is amenable.
-Hypotheses: Each dalpist is a rompus. Dalpists are large. Edward is a dalpist. Pamela is a felper.
-Multiple hypothesis; height-4 ontology tree.
+Hypotheses: Each dalpist is a rompus.
+Dalpists are large.
+Edward is a dalpist.
+Pamela is a felper.
+Multiple hypothesis example; height-4 ontology tree.
 World Model: Every kurpor is a twimpee.
@@
 Mark is translucent.
-Hypotheses: Dalpists are rompuses. Every dalpist is liquid. Emily is a gergit. Sarpers are large. Timples are salty. William is a rimpee. Tumpuses are brown. Orgits are translucent
-Appendix C Algorithms
-Refer to Algorithm 1 and 2.
-if task is infer property then
- select a member of the root node
- randomly select a member of the leaf nodes
- randomly select a member of nonleaf nodes
- generates the observations where selected members have the property
-end if
-if task is infer membership then
- select a random leaf node
- randomly select a property of the leaf node
- randomly select a property of the root node
- randomly select a property of the selected node’s non-root ancestors.
- generates the observations where this entity have all properties
-end if
-if task is infer subtype relation then
- select a member of the root node
- randomly select a member of the leaf nodes
- randomly select a member of nonleaf nodes
- generates the observations where selected members also belong to a supertype
-end if
-return all world model axioms, observations, and the ground truth hypothesis
-Algorithm 1 Generate a single hypothesis
-// First, build the ontology tree
-// $Node()$ will initialize a node object
-// Add a root node
-$tree\_nodes\leftarrow[[Node()]]$
-for $i$ from 1 to h1h-1 do
-$current\_nodes\leftarrow[]$
-for $parrent\_node\in tree\_nodes[-1]$ do
-  $u\leftarrow\texttt{random}(0,1)$
-  if $u>0.5$ then
-   for $j$ from 1 to 2 do
-    current_nodes.append(Node())current\_nodes.append(Node())
-   end for
-  else
-   for $j$ from 1 to 3 do
-    current_nodes.append(Node())current\_nodes.append(Node())
-   end for
-  end if
-end for
-tree_nodes.append(current_nodes)tree\_nodes.append(current\_nodes)
-end for
-for $current\_nodes\in tree\_nodes$ do
-for $node\in current\_nodes$ do
-  // Allocate three members and properties
-  for $j$ from 1 to 3 do
-   $node.members.append(RandomMember())$
-   $node.properties.append(RandomProperty())$
-  end for
-end for
-end for
-if single hypothesis then
- Randomly select a task type
- generates a single hypothesis
-else
- generate 1 to 3 hypothesis per tree layer
-1end if
-Algorithm 2 Generate a reason example
-Appendix D More results
+Hypotheses: Dalpists are rompuses.
+Every dalpist is liquid.
+Emily is a gergit.
+Sarpers are large.
+Timples are salty.
+William is a rimpee.
+Tumpuses are brown.
+Orgits are translucent.
+Appendix D Details of GPT-4o paraphrase
+D.1 Prompts
+We present our prompts for GPT-4o to paraphrase synthetic reasoning questions into more natural ones in Figure 6.
+Prompt for Question Rewriting Instruction We prompt GPT-4o to rewrite synthetic reasoning questions into more natural, human-readable language while preserving the original semantics and reasoning structure.
+Prompt Template You are a professional editor specializing in rewriting synthetic reasoning questions into natural, human-readable language.
+Rewrite the following question to be clear, fluent, and natural while preserving all logical structure and meaning.
+Do not simplify the reasoning steps, remove constraints, or change the problem’s difficulty.
+Return only the improved question with the exact same semantics, but in more natural language.
+Input question: {{QUESTION}}
+Figure 6: Prompt used to paraphrase synthetic reasoning questions into more natural language.
+D.2 Manual inspection
+We also manually verify that the generated questions from GPT-4o are sound by randomly sampling 200 reasoning questions of varying difficulty and manually checking that GPT-4o’s output does not alter the problem’s semantics.
+We conduct this study and iteratively improve the prompt until achieving 100% accuracy.
+Appendix E Details of hypothesis quality metric
+E.1 Theoretical justification
+Our quality metric can be motivated by an argument from Bayesian decision theory.
+Let $T$ be a theory which consists of a set of observations $o_{1},\ldots,o_{n}$ and a set of proofs $\pi_{1},\ldots,\pi_{n}$ where each proof $\pi_{i}$ explains the observation $o_{i}$ (i.e., the conclusion of $\pi_{i}$ is $o_{i}$).
+Let
+$p(T)=\prod_{i=1}^{n}p(\pi_{i}\mid\pi_{1},\ldots,\pi_{i-1})\,p(o_{i}\mid\pi_{i})$
+be a prior distribution over the theory. Ideally, during inductive/abductive reasoning, given a new observation $o_{n+1}$, we aim to find the proof $\pi_{n+1}$ that maximizes the posterior
+$p(\pi_{n+1}\mid\pi_{1},\ldots,\pi_{n},o)$ $\propto p(\pi_{n+1}\mid\pi_{1},\ldots,\pi_{n})$
+$\quad\cdot p(o_{n+1}\mid\pi_{n+1})$
+Since
+$p(o_{i}\mid\pi_{i})=\mathds{1}\{o_{i}\text{ is the conclusion of }\pi_{i}\},$
+the posterior is equal to
+$p(\pi_{n+1}\mid\pi_{1},\ldots,\pi_{n})$
+whenever the conclusion of $\pi_{n+1}$ is $o_{n+1}$.
+Now suppose we make the mild assumption that $p(T)$ is such that $p(T)>p(T^{\prime})$ whenever $T$ has fewer axioms in its proofs than does $T^{\prime}$.
+Therefore, $p(T)$ is maximized if $\pi_{n+1}$ minimally introduces new axioms (and equivalently, maximally reuses existing axioms).
+Note that the quality metric $g(\pi_{n+1})$ is the ratio of two quantities: (1) the average number of times each premise $h\in\pi_{n+1}$ appears in $\pi_{1},\ldots,\pi_{n}$, and (2) the average number of times each premise in the ground-truth proof appears in $\pi_{1},\ldots,\pi_{n}$.
+We specifically selected each ground-truth proof to be the unique proof with the minimum number of new axioms.
+If we consider two proofs $\pi_{n+1}$ and $\pi^{\prime}_{n+1}$ such that $\pi_{n+1}$ has more new axioms than $\pi^{\prime}_{n+1}$, then the sum in the numerator of $g(\pi_{n+1})$ would contain more zero terms than that of $g(\pi^{\prime}_{n+1})$, and so
+$g(\pi_{n+1})<g(\pi^{\prime}_{n+1})$
+necessarily.
+Thus, with mild assumptions on the prior over the theory $p(T)$, our proposed quality metric provides an easy-to-compute proxy for the posterior probability of the theory, where more probable theories will have higher quality scores.
+Furthermore, our quality metric is agnostic to the specific choice of the prior $p(T)$.
+E.2 Human validation experiments
+To evaluate whether the proposed metric aligns with human judgments of hypothesis quality, we conduct a human evaluation study.
+We recruit three independent annotators with graduate-level STEM degrees in US R1 universities.
+The evaluation dataset contains 100 reasoning questions, each associated with three candidate hypotheses.
+For each question, annotators are asked to select the hypothesis that better explains observations and is most concise.
+The presentation order of hypotheses is randomized to avoid positional bias.
+Each reasoning question is evaluated by three annotators, and the final preference label is determined by majority vote.
+To quantify alignment with human judgments, we measure agreement accuracy, defined as the percentage when the proposed metric assigns the highest score to the human-preferred hypothesis.
+The proposed metric achieves an agreement of 79%, substantially outperforming a random baseline (33.3%) and a length-based heuristic baseline (51%).
+Human annotations exhibit substantial agreement (Fleiss’ $\kappa$ = 0.75), indicating that the evaluation task is reasonably well-defined and that the proposed metric captures meaningful aspects of hypothesis quality.
+Please refer to Figure 7 for the instruction to annotators.
+Instruction for annotators --- You will be given a set of reasoning questions.
+Each question includes: * A description of observations * Three candidate hypotheses that attempt to explain these observations --- Your task is to select the single best hypothesis according to the following criteria: * The hypothesis should account for the given observations coherently and logically.
+* The hypothesis should be as concise as possible.
+--- Important Guidelines * Select only one hypothesis per question.
+* Do not choose based on writing style, fluency, or formatting alone.
+* Focus on the validity and quality of hypotheses, not superficial features.
+* The order of hypotheses is randomized; treat each option independently.
+* If multiple hypotheses seem similar, choose the one that is most concise.
+--- Your selection will be used to determine the hypothesis that best explains the observations.
+Figure 7: Instruction for annotators.
+Appendix F Algorithms
+Algorithms 1 and 2 describe how we generate ontology trees and hypotheses for the examples in our framework.
+1function generate_hypothesis(ontology tree $T$)
+2    if task is “infer property”
+3       randomly select a member of the root node of $T$
+4       randomly select a member of a leaf node of $T$
+5       generate the observations where selected members have the property
+6
+7   else if task is “infer membership”
+8       select a random leaf node of $T$
+9       randomly select a property of the leaf node
+10       randomly select a property of the root node
+11       randomly select a property of the selected node’s non-root ancestors
+12       generate observations where this entity has all properties
+13
+14   else if task is “infer subtype relation”
+15       randomly select a member of the root node of $T$
+16       randomly select a member of a leaf node of $T$
+17       randomly select a member of a non-leaf of $T$
+18       generate the observations where selected members also belong to a supertype
+19
+20   return all world model axioms, observations, and the ground-truth hypothesis
+Algorithm 1 Generate a single hypothesis.
+1function generate_example()
+2    /* first, build the ontology tree */
+3    /* node() initializes an empty node object */
+4    /* add a root node */
+5    $tree\_nodes\leftarrow$ [[node()]]
+6    for $i=1,\ldots,h-1$ do
+7       $current\_nodes\leftarrow[]$
+8       for $parent\_node\in tree\_nodes[-1]$ do
+9          $u\leftarrow$ random(0,1)
+10          if $u>\frac{1}{2}$
+11             repeat 2 times
+12                current_nodescurrent\_nodes.append(node())
+13
+14
+15         else
+16             repeat 3 times
+17                current_nodescurrent\_nodes.append(node())
+18
+19
+20
+21      tree_nodestree\_nodes.append(current_nodescurrent\_nodes)
+22
+23   for $current\_nodes\in tree\_nodes$ do
+24       for $node\in current\_nodes$ do
+25          /* allocate three members and properties */
+26          repeat 3 times
+27             nodenode.members .append( random_member())
+28             nodenode.properties .append( random_property())
+29
+30
+31
+32   if this is a single-hypothesis example
+33       randomly select a task type
+34       generate_hypothesis(tree_nodestree\_nodes)
+35
+36   else generate 1 to 3 hypotheses per tree layer
+37
+Algorithm 2 Generate an abductive/inductive reasoning example.
+Appendix G Results with absolute accuracy
 Here we list the absolute results of Figure 4 in Figure 8.
@@
 Figure 8: Absolute results of weak accuracy, strong accuracy, and quality versus the height of the ontology tree with multiple hypotheses under zero-shot (top), 8-shot in-distribution (middle), and 8-shot out-of-distribution (bottom).
-Appendix E LLM example responses
-Self-verification
-World Model: Amy is a dalpist. Pamela is a dalpist. Jerry is a dalpist.
-Observations: Amy is rainy. Jerry is rainy. Pamela is rainy.
-Hypotheses: Dalpists are rainy.
-LLM Output: <think> … Suppose Dalpists are rainy. Since Amy and Jerry and Pamela are dalpists, they are all rainy. This hypothesis can explain all observations. </think>
-Wrong ontology direction
-World Model: Porpors are dalpists.
+Appendix H Error analysis
+H.1 Common error types and example responses
+In this section, we detail the most common types of errors that LLMs produce in our experiments.
+For each of the error types listed below, we provide an example consisting of a world model, list of observations, ground-truth hypotheses, and the mistaken LLM output.
+Error Type 1: Wrong ontology direction World Model: Porpors are dalpists.
 Nicole is a gergit.
@@
 Nicole is moderate.
-Ground truth hypotheses: Dalpists are rompuses. Every dalpist is liquid. Each gergit is moderate. Barbara is a porpor.
-LLM Output: … Rompuses are dalpists. …
-Ignore the ontology and produce unnecessary hypotheses
-World Model: David is a sorple.
+Ground-truth hypotheses: Dalpists are rompuses.
+Every dalpist is liquid.
+Each gergit is moderate.
+Barbara is a porpor.
+Mistaken LLM Output: … Rompuses are dalpists.
+…
+Error Type 2: Ignore the ontology and produce unnecessary hypotheses World Model: David is a sorple.
 Each sorple is a dropant.
@@
 Every scrompist is a gomper.
-Observations: David is rainy. Brian is rainy. Angela is rainy.
-Ground truth hypotheses: Every gomper is rainy.
-LLM Output: … Every gomper is rainy. Every dropant is rainy. …
-Fall back to rival hypothesee
-World Model: Rorpants are dolpers.
+Observations: David is rainy.
+Brian is rainy.
+Angela is rainy.
+Ground-truth hypotheses: Every gomper is rainy.
+Mistaken LLM Output: … Every gomper is rainy.
+Every dropant is rainy.
+…
+Error Type 3: Fall back to rival hypotheses World Model: Rorpants are dolpers.
 Every worple is a wumpus.
@@
 Michael is a gwompant.
-Observations: Shirley is a borpin. Christopher is pale. Emma is a borpin. Brian is salty. Michael is salty. Christopher is a borpin. Christopher is cold. Ashley is salty. Christopher is moderate.
-Ground truth hypotheses: Each gwompant is a borpin. Each gwompant is salty. Christopher is a drompee.
-LLM Output: … Shirley is a borpin. …
-Hallucinated entities
-World Model: Jonathan is a rifpist.
+Observations: Shirley is a borpin.
+Christopher is pale.
+Emma is a borpin.
+Brian is salty.
+Michael is salty.
+Christopher is a borpin.
+Christopher is cold.
+Ashley is salty.
+Christopher is moderate.
+Ground-truth hypotheses: Each gwompant is a borpin.
+Each gwompant is salty.
+Christopher is a drompee.
+Mistaken LLM Output: … Shirley is a borpin.
+…
+Error Type 4: Hallucinated entities World Model: Jonathan is a rifpist.
 Yimples are dull.
@@
 Each yimple is a rifpist.
-Observations: Mary is not overcast. Scott is not overcast. Jonathan is not overcast. Samantha is a kurpor. Mary is aggressive. Jessica is a kurpor. Olivia is a kurpor. Mary is small. Mary is dull.
-Ground truth hypotheses: Every jelgit is a kurpor. Jelgits are not overcast. Mary is a yimple.
-LLM Output: … Mary is a dolphant. …
-Confuse members with concepts
-World Model: Delpees are impuses.
-Yumpuses are delpees.
-Rebecca is an impus.
-Kathleen is a rifpist.
-Each impus is rainy.
-Thomas is a delpee.
-Lompees are arpers.
-Each bempin is an impus.
-Stephen is a bempin.
-Every impus is shy.
-Laura is a yumpus.
-Every arper is an impus.
-Stephanie is an arper.
-Lompees are moderate.
-Rifpists are bempins.
-Observations: Frank is moderate. Frank is rainy. Frank is shy.
-Ground truth hypotheses: Frank is a lompee.
-LLM Output: … Lompee is Frank. …
+Observations: Mary is not overcast.
+Scott is not overcast.
+Jonathan is not overcast.
+Samantha is a kurpor.
+Mary is aggressive.
+Jessica is a kurpor.
+Olivia is a kurpor.
+Mary is small.
+Mary is dull.
+Ground-truth hypotheses: Every jelgit is a kurpor.
+Jelgits are not overcast.
+Mary is a yimple.
+Mistaken LLM Output: … Mary is a dolphant.
+…
+H.2 Details of error analysis
+We first randomly sampled 200 LLM example responses to reasoning questions of varying difficulty.
+Then, we manually inspected them to identify recurring failure patterns as summarized in the paper.
+To further quantitatively study the error patterns of all responses, we use an LLM-as-a-judge (llmasjudge) and prompt GPT-4o with in-context examples.
+We use the following prompt in Figure 10 for the LLM-as-a-judge.
+We categorize the error pattern distribution by ontology tree height and present the results in Figure 9.
+As shown in the figure, Error Type 2 (Ignore the ontology and produce unnecessary hypotheses) is the most significant error across all ontology tree heights, indicating that LLMs can generate weakly correct hypotheses but fail to consider the given ontology tree and optimize them in the sense of Occam’s razor.
+Figure 9: Distribution of error types across ontology tree heights.
+Prompt for Error Analysis Prompt Template You are an expert evaluator of reasoning quality.
+Your task is to analyze a model’s reasoning process for a question and determine the type of error it contains.
+## Task Given: * A question
+* The model’s reasoning process
+* The correct answer
+Identify which error type best describes the mistake in the reasoning.
+Choose exactly one error type from the following categories.
+--- ## Error Types ** Error Type 1: Wrong ontology direction** The hypothesis contains the wrong ontology direction.
+** Error Type 2: Fall back to trivial hypotheses** The hypothesis is the observation itself.
+** Error Type 3: Ignore the ontology and produce unnecessary hypotheses** The hypothesis is technically correct but unnecessary given the ontology.
+** Error Type 4: Hallucinated entities** The reasoning relies on non-existent entities.
+--- ## Examples ### Example 1 (Wrong ontology direction) <omitted for space, see §H.1 for examples> --- ### Example 2 (Fall back to trivial hypotheses) <omitted for space, see §H.1 for examples> --- ### Example 3 (Ignore the ontology and produce unnecessary hypotheses) <omitted for space, see §H.1 for examples> --- ### Example 4 (Hallucinated entities) <omitted for space, see §H.1 for examples> --- ## Now evaluate the following case Question: {{QUESTION}} Model Reasoning: {{MODEL_REASONING}} Correct Answer: {{CORRECT_ANSWER}} --- ## Output Format Error Type: <one of: Wrong ontology direction | Fall back to trivial hypotheses | Ignore the ontology and produce unnecessary hypotheses | Hallucinated entities>
+Figure 10: Prompt used to analyze common error patterns.